### v1

# Flexible Deep Neural Networks for Partially Linear Survival Data

Abstract
We propose a flexible deep neural network (DNN) framework for modeling survival data within a partially linear regression structure. The approach preserves interpretability through a parametric linear component for covariates of primary interest, while a nonparametric DNN component captures complex time–covariate interactions among nuisance variables. We refer to the method as FLEXI–Haz, a FLEXIble Hazard model with a partially linear structure. In contrast to existing DNN approaches for partially linear Cox models, FLEXI–Haz does not rely on the proportional hazards assumption. We establish theoretical guarantees: the neural network component attains minimax-optimal convergence rates that depends on composite Hölder classes, and the linear estimator is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Extensive simulations and real-data analyses demonstrate that FLEXI–Haz provides accurate estimation of the linear effect, offering a principled and interpretable alternative to modern methods based on proportional hazards. Code for implementing FLEXI-Haz, as well as scripts for reproducing data analyses and simulations is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
keywords: Deep neural networks; Interpretability; Partial linear regression; Survival analysis; Time to event data.
1 Introduction
Deep neural networks (DNNs) are widely used in modern data analysis because of their ability to capture complex nonlinear patterns, handle high-dimensional inputs, and scale to large datasets. They have achieved strong performance in diverse domains such as image recognition, speech processing, and natural language understanding, often improving on traditional statistical and machine learning methods (lecun2015deep). Continued progress in architectures, optimization, and computational resources has further expanded their range of applications in both scientific and industrial domains.
Deep neural networks for survival analysis.
Survival analysis aims to model the distribution of event times in the presence of censoring, with applications in medicine, public health, epidemiology, engineering, and finance (kalbfleisch2011; klein2006survival). In recent years, DNNs have become increasingly popular in survival analysis thanks to their ability to model complex nonlinear relationships between covariates and event times and to scale efficiently to high-dimensional data. Many existing approaches extend the classical semi-parametric Cox proportional hazards (PH) model (cox1972regression) by replacing the parametric linear risk score with a neural network, while using the negative partial likelihood as the network’s loss function. Early work by faraggi1995neural employed a single hidden-layer network in place of the linear component, and katzman2018deepsurv extended this to deeper architectures (DeepSurv), demonstrating improved concordance in several benchmark datasets. Variants of this approach have been proposed for genomics (yousefi2017predicting; ching2018coxnnet), medical imaging (haarburger2019image; li2019deep), and recurrent or discrete-time formulations (giunchiglia2018rnnsurv; gensheimer2019nnet). While these PH-based DNN methods can yield strong empirical performance, they inherit the PH restriction, which may limit their ability to capture rich time-dependent effects.
Non-PH modeling with deep learning.
To address the limitations of the PH assumption, a growing body of work has focused on DNN-based models that allow covariate effects to vary with time. One strategy augments the network input with time itself, enabling non PH while retaining an unspecified baseline hazard, as in Cox-Time of kvamme2019time. Other approaches abandon the Cox structure entirely in favor of modeling the event time distribution or hazard function directly. In particular, hu2023conditional proposed a framework for estimating the full survival function, using a likelihood function derived from the continuous-time hazard without introducing a baseline hazard function. Their formulation naturally accommodates time-varying covariates through a data-expansion scheme and applies optimization to the log-hazard. By sidestepping the baseline hazard and proportionality constraints, such approaches offer greater modeling flexibility and can better reflect complex temporal dynamics in the hazard. The utility of this approach is illustrated through simulations and real-data analyses, but no asymptotic results are provided.
Semiparametric models and deep learning.
Alongside fully nonparametric formulations, there is considerable interest in semiparametric survival models that combine a parametric component for covariates of primary interest with a flexible, nonparametric component for high-dimensional or nuisance covariates. This structure preserves interpretability and enables formal inference for the parametric component, while leveraging modern machine learning to estimate the nuisance function. Theoretical foundations for this paradigm have been established in the non-survival setting by schmidthieber2020nonparametric, who derived minimax-optimal rates for ReLU-activated DNNs in nonparametric regression, and extended to survival analysis by zhong2022deep in the partially linear Cox model. In the latter, the log-hazard at time $t$ is modeled as $\log h_{0}(t)+\mbox{${\zeta}$}^{\top}{\bf Z}+f({\bf X})$, where $h_{0}(t)$ is an unspecified baseline hazard function, ${\bf Z}\in\mathbb{R}^{p}$ is the vector of covariates of primary interest, ${\bf X}\in\mathbb{R}^{d}$ is the vectors of nuisance covariates, and $\mbox{${\zeta}$}\in\mathbb{R}^{p}$ is the main parameter of interest, estimated at the semiparametric efficiency bound via partial likelihood. The nuisance function $f$ is approximated by a DNN. This approach achieves both optimal convergence for $g$ and $\sqrt{n}$-consistency for $\zeta$} under the PH assumption. The work of zhong2022deep demonstrates how semiparametric theory and deep learning can be integrated to yield interpretable, efficient estimators in survival analysis, but like other Cox-based methods, it remains tied to PH and baseline hazard estimation.
Contributions.
In this work, we develop a semiparametric survival model that extends the partially linear framework of zhong2022deep to the non-PH setting. We model the log-hazard function at time $t$ as $\mbox{${\theta}$}_{o}^{\sf T}{\bf Z}+g_{o}(t,{\bf X})$, where $\mbox{${\theta}$}_{o}\in\mathbb{R}^{p}$ is the parameter of primary interest and $g_{o}$ is an unknown nuisance function approximated by a DNN. This formulation allows arbitrary interactions between the nuisance covariates $\bf X$} and time $t$, naturally accommodates time-dependent effects and time-dependent covariates within $\bf X$}, and does not require specifying or estimating a baseline hazard function.
A central methodological difference from existing semiparametric approaches is that we work with the full likelihood, rather than the partial likelihood commonly used in Cox-type models. Despite removing the PH structure, we establish that the estimator of the parametric component $\mbox{${\theta}$}_{o}$ remains $\sqrt{n}$-consistent, asymptotically normal, and semiparametric efficient. In particular, we show that the theoretical guarantees of zhong2022deep, developed for the partially linear Cox model, continue to hold in the substantially more general non-PH setting, provided the hazard is almost surely bounded away from zero.
While parts of our analysis build on the approximation theory of schmidthieber2020nonparametric and the high-level strategy of zhong2022deep, several key components of our asymptotic theory are fundamentally different. The core technical challenges arise from the fact that our neural network estimator is a function of both $t$ and $\bf X$}, rather than $\bf X$} alone, and the fact that we use the full likelihood rather than a profiling argument. These differences require new empirical process tools, a new characterization of the nuisance tangent space, and a new argument for controlling the bias induced by the DNN approximation.
Finally, when the linear effect $\mbox{${\theta}$}_{o}^{\top}{\bf Z}$ is removed and $\bf Z$} is absorbed into $g_{o}$, our model collapses to the conditional hazard model of hu2023conditional. Our analysis therefore fills an important theoretical gap by providing the first asymptotic guarantees for that model as well.
Code for implementing FLEXI-Haz, as well as scripts for reproducing all data analyses and simulations, is available at GitHub site https://github.com/AsafBanana/FLEXI-Haz.
2 Model Formulation
We consider right-censored survival data, where for each individual the true event time is denoted by $U$ and the censoring time by $C$. We observe only $T=\min(U,C)$ and $\Delta=\mathbf{1}(U\leq C)$, indicating the observed time and whether the event occurred. Let the covariate vector be $({{\bf Z}}^{\sf T},{{\bf X}}^{\sf T})^{\sf T}$, where ${{\bf Z}}\in\mathbb{R}^{p}$ contains the variables of primary scientific interest and ${{\bf X}}\in\mathbb{R}^{d}$ denotes nuisance covariates. Our goal is to characterize the effect of $\bf Z$} on the event time while appropriately adjusting for $\bf X$}. For example, $\mathbf{X}$ may represent a high-dimensional set of genomic or imaging features, whereas $\mathbf{Z}$ may contain key clinical variables such as treatment assignment or targeted biomarkers. Understanding the effect of $\bf Z$} while properly adjusting for $\bf X$} is essential for identifying clinically actionable signals. For simplicity, we focus here on time-independent covariates; the extension to time-dependent covariates will be discussed in Section 7. We assume that $U$ and $C$ are conditionally independent given $(\mathbf{X},\mathbf{Z})$. The observed data consist of $n$ independent replicates, $\{(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$.
Under the above setting, the Cox PH model (cox1972regression) specifies the instantaneous hazard function at time $t>0$ as
$h_{1}(t\mid{{\bf X}},{{\bf Z}})=h_{0}(t)\exp({\mbox{${\alpha}$}}_{1}^{\top}{{\bf X}}+{\mbox{${\alpha}$}}_{2}^{\top}{{\bf Z}})\,$,
where $h_{0}$ is an unspecified baseline hazard function and ${\mbox{${\alpha}$}}_{1}\in\mathbb{R}^{d}$ and ${\mbox{${\alpha}$}}_{2}\in\mathbb{R}^{p}$ are unknown parameters. The model remains widely used due to its interpretability, semiparametric structure, and the availability of efficient estimation procedures. DeepSurv (katzman2018deepsurv) extends the Cox model by replacing the linear risk score $\exp({\mbox{${\alpha}$}}_{1}^{\top}\mathbf{X}+{\mbox{${\alpha}$}}_{2}^{\top}\mathbf{Z})$ with a flexible nonlinear function $f_{1}({{\bf X}},{{\bf Z}})$ estimated by a DNN. Cox-Time (kvamme2019time) further generalized this formulation by replacing $f_{1}$ with a time-depend function, $f_{2}(t,{{\bf X}},{{\bf Z}})$, thereby relaxing the PH assumption and permitting time-covariate interactions. Because $f_{2}$ is time-dependent while the baseline hazard is left unspecified, identifiability becomes nontrivial; kvamme2019time address this challenge through a numerical normalization strategy. More recently, hu2023conditional proposed a baseline-free hazard model of the form
$h_{3}(t\mid{{\bf X}},{{\bf Z}})=\exp\{f_{3}(t,{{\bf X}},{{\bf Z}})\}\,$,
which provides full flexibility for modeling non-proportional, time-varying effects via a DNN. Their estimation procedure is evaluated only through simulations, with no accompanying asymptotic guarantees. Although these DNN-based approaches are highly flexible, they do not provide interpretable estimates for the effect of $\bf Z$}}. To address this limitation, zhong2022deep introduced the deep partially linear Cox model
$h_{4}(t\mid\mathbf{X},\mathbf{Z})=h_{0}(t)\exp\{\mbox{${\zeta}$}^{\top}\mathbf{Z}+f(\mathbf{X})\}\,$,
which augments the Cox framework by separating time component from the covariates and retaining a parametric term $\mbox{${\zeta}$}^{\top}\mathbf{Z}$ for interpretability, while modeling nuisance effects through a DNN. This structure enables valid inference for $\zeta$} while still allowing complex, nonlinear effect of $\bf X$}}.
In this work, we extend the deep partially linear framework of zhong2022deep to the non-PH setting by considering a baseline-free model of the form
$h_{o}(t\mid\mathbf{X},{{\bf Z}})=\exp\{\mbox{${\theta}$}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})\},$ (1)
where $g_{o}(t,\mathbf{X})$ is estimated by a DNN. Our formulation retains the interpretability and inference guarantees for the parametric component ${\mbox{${\theta}$}}_{o}$, but generalizes the nuisance component from $f(\mathbf{X})$ to $g_{o}(t,\mathbf{X})$, thereby allowing fully flexible, time-varying effects and removing the reliance on a baseline hazard function. This extension overcomes the PH restriction inherent in zhong2022deep, resolves the non-identifiability issues that arise when time is included in the DNN, and greatly expands the class of survival models that can be handled within a semiparametric, partially linear framework.
As we demonstrate in Section 6, restricting the nuisance function to depend only on only on $\mathbf{X}$, as in the partially linear Cox model, can induce substantial bias in the estimation of the linear effect whenever the nuisance covariates exhibit time-varying effects. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ eliminates this source of bias but introduces significant methodological challenges. First, the negative partial likelihood can no longer be used as the DNN loss function. Instead, the full-likelihood is adopted which includes the predictable cumulative hazard that must be numerically approximated at every training step, increasing the computational burden. Second, the presence of this integral markedly complicates the asymptotic analysis: the time dimension introduces additional stochastic terms that must be controlled, requiring new empirical process arguments beyond those used for time-invariant nuisance functions. Our estimation framework in Section 3 addresses these computational difficulties, and the technical developments leading to the asymptotic results are presented in Section 4.3.
2.1 Composite Hölder Class of $g_{o}$
To formalize the complexity assumptions on the nuisance component $g_{o}$, we follow the compositional function framework of juditsky2009 and schmidthieber2020nonparametric, and use the notation of zhong2022deep. We assume that the nuisance function $g_{o}$ admits a hierarchical, low-dimensional representation of the form
$g_{o}=g_{q}\circ g_{q-1}\circ\cdots\circ g_{1}\circ g_{0},$
with the convention that $(g\circ f)(x)=g(f(x))$. This composition is associate with a dimension vector ${\bf d}^{*}=(d^{*}_{0},\ldots,d^{*}_{q+1})\in\mathbb{N}_{+}^{q+2}$ and an intrinsic dimension vector $\tilde{\bf d}=(\tilde{d}_{0},\ldots,\tilde{d}_{q})\in\mathbb{N}_{+}^{q+1}$, $\tilde{d}_{\ell}\leq d^{*}_{\ell}$ for each $\ell$. The outermost input is $d^{*}_{0}=r=1+d$, corresponding to observed times and covariates $\mathbf{X}$, and the final output is scalar $d^{*}_{q+1}=1$. For each layer $l$, $l=0,\ldots,q$, the vector-valued mapping $g_{\ell}=(g_{\ell 1},\ldots,g_{\ell\,d^{*}_{\ell+1}})^{\top}$ maps the bounded set $a_{\ell},b_{\ell}]^{\,d^{*}_{\ell}}$ into $a_{\ell+1},b_{\ell+1}]^{\,d^{*}_{\ell+1}}$, where $a_{l},b_{l}\in\mathbb{R}$ are fixed finite constants ensuring each intermediate representation remains bounded. Additionally, each component $g_{\ell j}$ depending on at most $\tilde{d}_{\ell}$ coordinates of its input.
For illustration, consider $f:\mathbb{R}^{6}\to\mathbb{R}$ defined by
$f(x_{1},\ldots,x_{6})\;=\;\sin(x_{1}+x_{2})\;+\;\log(1+x_{3})\;+\;x_{4}^{2}.$
which depends on low-dimensional groupings of the input coordinates. A valid three-layer decomposition is $g\;=\;g_{2}\circ g_{1}\circ g_{0}$ where
$g_{0}:\mathbb{R}^{6}\to\mathbb{R}^{3}\,,\quad g_{0}(x)\;=\;\big(\,x_{1}+x_{2},\;\;x_{3},\;\;x_{4}\,\big)\,$,
$g_{1}:\mathbb{R}^{3}\to\mathbb{R}^{3}\,,\quad g_{1}(y_{1},y_{2},y_{3})\;=\;\big(\,\sin(y_{1}),\;\;\log(1+y_{2}),\;\;y_{3}^{2}\,\big)\,$,
$g_{2}:\mathbb{R}^{3}\to\mathbb{R}\,,\quad g_{2}(z_{1},z_{2},z_{3})\;=\;z_{1}+z_{2}+z_{3}\,.$
For this representation, ${\bf d}^{*}=({d^{*}_{0}},{d^{*}_{1}},{d^{*}_{2}},{d^{*}_{3}})=(6,3,3,1)$, and the intrinsic dimension vector is $\tilde{\bf d}=(\tilde{d}_{0},\tilde{d}_{1},\tilde{d}_{2})=(2,1,3)$. That is, components of $g_{0}$ depend on at most two inputs, those of $g_{1}$ are univariate, and $g_{2}$ combines all three outputs. Other valid decompositions may exist. In general, convergence rates depend on the choice of the optimal decomposition used in the asymptotic analysis.
We also assume that each component function satisfies a standard Hölder–smoothness condition. Let $\alpha>0$ and $M>0$. A function is said to have Hölder smoothness index $\alpha$ if all mixed partial derivatives up to order $\lfloor\alpha\rfloor$ exist and are uniformly bounded, and if the derivatives of order exactly $\lfloor\alpha\rfloor$ are $(\alpha-\lfloor\alpha\rfloor)$–Hölder continuous. Here, $\lfloor\alpha\rfloor$ denotes the greatest integer strictly smaller than $\alpha$. The ball of $\alpha$-Hölder functions with radius $M$ consists of all such functions whose derivatives and Hölder constants are bounded by $M$. Formally, for a domain $\mathbb{D}\subset\mathbb{R}^{m}$, the Hölder ball is defined as
$\mathcal{H}^{\alpha}_{m}(\mathbb{D},M)=\Big\{g:\mathbb{D}\to\mathbb{R}:\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|<\alpha}\|\partial^{\boldsymbol{\beta}}g\|_{\infty}+\sum_{\boldsymbol{\beta}:|\boldsymbol{\beta}|=\lfloor\alpha\rfloor}\sup_{u,v\in\mathbb{D},u\neq v}\frac{|\partial^{\boldsymbol{\beta}}g(u)-\partial^{\boldsymbol{\beta}}g(v)|}{\|u-v\|^{\alpha-\lfloor\alpha\rfloor}}\leq M\Big\},$
where $\boldsymbol{\beta}^{\sf T}=(\beta_{1},\ldots,\beta_{m})\in\mathbb{N}^{m}$ is a multi-index $\partial^{\boldsymbol{\beta}}=\partial^{\beta_{1}}\cdots\partial^{\beta_{m}}$, $|\boldsymbol{\beta}|=\sum_{j=1}^{m}\beta_{j}$, and $\|g\|_{\infty}$ is the sup–norm of a function over its domain.
To formalize the function class used for the nuisance component, let $\boldsymbol{\alpha}^{\sf T}=(\alpha_{0},\ldots,\alpha_{q})\in\mathbb{R}_{+}^{q+1}$ denote the layerwise smoothness parameters, $\bf d$} the dimension vector, and $\tilde{{\bf d}}$ the intrinsic dimension vector. The composite Hölder class is defined as
$\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)=\Big\{g=g_{q}\circ\cdots\circ g_{0}:g_{\ell j}\in\mathcal{H}^{\alpha_{\ell}}_{\tilde{d}_{\ell}}([a_{\ell},b_{\ell}]^{\tilde{d}_{\ell}},M),\ |a_{\ell}|,|b_{\ell}|\leq M\Big\}.$
This class accomodates a broad range of structured functions, such as additive models, single- and multiple-index models, and hierarchical interactions, through the intrinsic dimensions $\tilde{d}_{\ell}$. Finally, we assume that the true nuisance function $g_{o}$ belongs to the composite class, $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
In the classical (non-composite) Hölder class $\mathcal{H}^{\alpha}_{d}$, the minimax-optimal rate is $n^{-\alpha/(2\alpha+d)}$ (tsybakov2009). For multilayer composite Hölder classes, schmidthieber2020nonparametric established matching minimax lower and upper bounds (up to polylogarithmic factors), showing that the optimal rate is governed by $\gamma_{n}\;=\;\max_{\ell=0,\ldots,q}\,n^{-\tilde{\alpha}_{\ell}/(2\tilde{\alpha}_{\ell}+\tilde{d}_{\ell})}$, where the effective smoothness indices are $\tilde{\alpha}_{\ell}\;=\;\alpha_{\ell}\prod_{k=\ell+1}^{q}(\alpha_{k}\wedge 1),\ \ a\wedge b=\min\{a,b\}$. Thus, each layer contributes to the overall difficulty of the estimation problem through its intrinsic dimension $\tilde{d}_{\ell}$ and effective smoothness $\tilde{\alpha}_{\ell}$. In Theorem 2 we show that our estimator achieves this optimal composite rate $\gamma_{n}$, up to polylogarithmic factors.
2.2 Neural Network Class
To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (schmidthieber2020nonparametric, and reference therein). This choice is motivated by two considerations: (i) ReLU networks are universal approximators capable of representing complex, high-dimensional functions; and (ii) their structure lends itself to sharp theoretical control through entropy bounds and sparsity constraints. We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r$ denotes the input dimension and $p_{K+1}=1$ the scalar output. Specifically, the network is defined by
$g(\mathbf{u})\;=\;\mathbf{W}_{K}\,\sigma\!\big(\mathbf{W}_{K-1}\,\sigma(\cdots\sigma(\mathbf{W}_{0}\mathbf{u}+\mathbf{v}_{0})\cdots)+\mathbf{v}_{K-1}\big)+\mathbf{v}_{K},$
where $\mathbf{u}\in\mathbb{R}^{r}$, $\sigma(x)=\max\{x,0\}$ is applied componentwise, $\mathbf{W}_{\ell}\in\mathbb{R}^{p_{\ell+1}\times p_{\ell}}$ are weight matrices, and $\mathbf{v}_{\ell}\in\mathbb{R}^{p_{\ell+1}}$ are bias vectors, $\ell=0,\ldots,K$.
For theoretical analysis we first restrict attention to networks with uniformly bounded parameters,
$\mathcal{G}(K,\boldsymbol{p})\;=\;\Big\{\,g:\ \max_{\ell=0,\ldots,K}\big(\|\mathbf{W}_{\ell}\|_{\infty}\vee\|\mathbf{v}_{\ell}\|_{\infty}\big)\leq 1\,\Big\}\,$,
which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms is adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$,. Following schmidthieber2020nonparametric we further impose sparsity. For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is define as
$\mathcal{G}(K,s,\boldsymbol{p},D)\;=\;\Big\{g\in\mathcal{G}(K,\boldsymbol{p})\ :\ \sum_{\ell=0}^{K}\big(\|\mathbf{W}_{\ell}\|_{0}+\|\mathbf{v}_{\ell}\|_{0}\big)\leq s,\ \ \|g\|_{\infty}\leq D\Big\},$
where $\|\cdot\|_{0}$ counts nonzero entries. Sparsity serves to control network complexity and is essential for obtaining optimal convergence rates. In the subsequent analysis we set the depth, width, and sparsity levels to balance bias-variance tradeoff. The bracketing entropy (vdvwellner1996, Chs. 2–3) of this class is then bounded using results of schmidthieber2020nonparametric, and Dudley’s entropy integral (vdvwellner1996, Thm. 2.14.1) is used to establish the required uniform empirical-process bounds.
3 FLEXI-Haz Procedure
Estimation is based on maximizing the full likelihood induced by the hazard model (1). Under noninformative and conditionally independent censoring given $({\bf X},{\bf Z})$, the corresponding full log–likelihood is proportional to
$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (2)
where $Y_{i}(t)\;=\;\mathbf{1}(T_{i}\geq t)$, $t\in[0,\tau]$, $i=1,\ldots,n$, are at-risk processes, and $\tau$ is a prespecified constant. Because $g$ is modeled by a DNN, the predictable cumulative hazard in (2) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization. We therefore introduce a grid $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ and approximate the integral by the Riemann–type sum
$\int_{0}^{\tau}Y_{i}(t)\,\exp\!\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\,dt\;\approx\;\sum_{j:\,Y_{i}(t_{j-1})=1}\big(\min\{T_{i},t_{j}\}-t_{j-1}\big)\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}.$
Defining $\Delta_{ij}=\mathbf{1}\{T_{i}\in(t_{j-1},t_{j}],\Delta_{i}=1\}$, we obtain the discretized likelihood
$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3)
In practice, discretizing the time axis on a fixed fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1. This construction automatically produces a dense grid in regions where failures occur, precisely where the integrand contributes most to the likelihood.
We consider two estimators. The oracle maximum–likelihood estimator is defined as
$(\widehat{}\mbox{${\theta}$},\widehat{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (4)
The estimator that is actually implemented replaces the at-risk integral with its discretized version $\widetilde{\ell}_{n}$. We therefore define the numerically integrated likelihood estimator as
$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (5)
Theorems 1–4 establish the asymptotic properties of (4), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (5) implemented in practice.
4 Asymptotic Analysis
4.1 Notation
Some additional notation will be used throughout. We write $a_{n}\lesssim b_{n}$ if there exists a constant $c>0$ such that $a_{n}\leq c\,b_{n}$ for all $n$, and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. For a column vector $\bf a$}, ${\bf a}^{\otimes 2}={\bf a}{\bf a}^{\sf T}$. Let $P_{0}$ denote the distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$. For any measurable function $h=h(T,{\bf X},{\bf Z})$ we define the unweighted $L^{2}_{P_{0}}$ norm $\|h\|^{2}_{L^{2}({P_{0}})}=\mathbb{E}_{0}\{h(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|h\|^{2}_{L^{2}(P_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,h(T,{\bf X},{\bf Z})^{2}\big\}$.
Let $\eta=(\mbox{${\theta}$},g)$, write
$\chi_{\eta}(t,{\bf X},{\bf Z})=g(t,{\bf X})\ +\ \theta^{\top}{\bf Z}\,$,
and denote the true parameter by $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$. For two parameters $\eta_{1},\eta_{2}$, we measure their respective event-weighted score distance by
$d(\eta_{1},\eta_{2})\ =\ \Big(\mathbb{E}_{0}\big[\Delta\big\{\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})\big\}^{2}\big]\Big)^{1/2}\ =\ \|\chi_{\eta_{1}}-\chi_{\eta_{2}}\|_{L^{2}({P_{0},\Delta})}\,.$
Finally, the population loss is defined as
$\ell_{0}(\eta)\;=\;\mathbb{E}_{0}\Big[\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\exp\!\big\{\chi_{\eta}(t,{\bf X},{\bf Z})\big\}\,dt\Big]\,.$ (6)
4.2 Assumptions
A1.
The failure time $U$ and censoring time $C$ are conditionally independent given $({\bf X}^{\top},{\bf Z}^{\top})$, and the censoring mechanism is noninformative in the sense that the censoring process is noninformative for $(\mbox{${\theta}$},g)$. The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
A2.
The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{p+q}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact. Without loss of generality, assume $\mathcal{X}\subset[0,1]^{d}$ and that $\|\mathbf{Z}\|_{2}\leq B_{{\bf Z}}$ almost surely. The joint density of $({\bf X},{\bf Z})$ is bounded away from zero on $\mathcal{X}\times\mathcal{Z}$. The parameter space for the finite-dimensional component is restricted to the compact set $\Theta\;=\;\{\vartheta\in\mathbb{R}^{p}:\ \|\vartheta\|\leq M_{\Theta}\}$. Finally, the study follows participants over a fixed, finite time horizon $\tau>0$.
A3.
The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$. In addition, the sieve is uniformly bounded, $\|g\|_{\infty}\leq D_{n}$ for a known sieve radius $D_{n}=O(1)$.
A4.
There exist $\delta_{1},\delta_{2}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$ and $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$.
A5.
The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term. Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{X}$ is positive definite, that is,
$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}^{\top}\right]\succ 0\,.$
Assumptions A1A3, A5 used also in zhong2022deep under the partial linear Cox regression model. A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (schmidthieber2020nonparametric). Assumption A2 is a standard assumption for semi- or non-parametric regression. Moreover, A1 and A2 imply that there exist constants $0<c<C<\infty$ (that depend on $D_{n},B_{{\bf Z}}$ and $M_{\Theta}$) such that for all $t\in(0,\tau]$ and $({\bf X},{\bf Z})\in\mathcal{X}\times\mathcal{Z}$, $c_{h}\leq h_{o}(t|{\bf X},{\bf Z})\leq C_{h}$. Assumption A5 is a standard idenfiability assumption in survival analysis. Assumption A4 is required since in the hazard function (1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disatengled, in contrast to the partial linear Cox model of zhong2022deep. Hence we need a non-vanishing chance to observe events over the entire study horizon for each covariate pattern.
4.3 Asymptotic Results
The proofs of all lemmas and theorems in this section appear in the appendix. We begin by establishing several basic properties of the model. A key step is to understand how the counting–process framework allows us to translate expressions involving at-risk process into simpler quantities expressed directly in terms of the event indicator $\Delta$. The next lemma formalizes this relationship, showing that these two representations are equivalent up to a multiplicative constant. This equivalence will be used repeatedly throughout the analysis.
Lemma 1.
Under Assumptions A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
$\mathbb{E}\!\left\{\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right\}\;\asymp\;\mathbb{E}\!\left\{\Delta\,f(T,{\bf X},{\bf Z})\right\}.$
The proof uses only the positivity of the hazard function together with the compensator identity for counting processes, and therefore applies uniformly to all functions that arise in our later analysis.
Lemma 2.
Let $\chi_{\eta_{j}}(t,{\bf X},{\bf Z})$, $j=1,2$, be bounded predictable processes satisfying
$\mathbb{E}\!\left\{\int_{0}^{\tau}\chi_{\eta_{j}}(t,{\bf X},{\bf Z})^{2}\,dt\right\}<\infty\qquad j=1,2\,.$
Define $\phi(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$. Under Assumptions A1, A2 and A4, there exist finite constants $0<c_{1}\leq c_{2}<\infty$ and $0<c_{3}\leq c_{4}<\infty$, depending only on $(\tau,\delta_{1},\delta_{2},c_{h},C_{h})$, such that
$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (7)
$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (8)
In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. and $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue-a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$, i.e. $\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes P_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincide almost everywhere. In addition, the lemma shows that the event-weighted norm and the corresponding unweighted norms are equivalent up to multiplicative constants. As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the usual $L^{2}$ norm and vice versa. This equivalence is used repeatedly throughout the theoretical development.
Next we study the shape of the population loss around $\eta_{o}=(g_{o},\mbox{${\theta}$}_{o})$.
Lemma 3.
Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2}.$
Lemma 3 shows that the population loss is well behaved in a neighbourhood of the truth: any deviation from $\eta_{o}$ incurs a loss of order $d(\eta,\eta_{o})^{2}$. Combined with Lemma 2, this implies that $\eta_{o}$ is the unique minimizer of the population loss over the square–integrable class under consideration. This population curvature is a key ingredient in the analysis that will later be paired with empirical process bounds to obtain convergence rates for $\widehat{g}$ and to derive the asymptotic distribution of $\widehat{\mbox{${\theta}$}}$.
We now move to the main asymptotic results. Our ultimate goal is to derive the asymptotic distribution of $\widehat{}\mbox{${\theta}$}$ and to show that it attains the semiparametric information bound. To this end, we first establish the convergence rate of the nuisance estimator $\widehat{g}$, show that this rate is essentially optimal, characterize the efficient score and information for $\mbox{${\theta}$}_{o}$, and then combine these ingredients to obtain an asymptotic distribution of $\widehat{}\mbox{${\theta}$}$. Finally, we verify that the numerical approximation used in practice does not alter the limit law.
We begin by quantifying how well the estimator $\widehat{g}$ recovers the unknown function $g_{o}$. The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{P_{0}}$ norm that naturally appears in the loss and in the distance function $d$. The rate is expressed in terms of the complexity parameter $\gamma_{n}$ and will later be used to control remainder terms in the expansion of the efficient score.
Theorem 1.
[…]
$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (9)
where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim P_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$. The quantity in (9) is random through its dependence on the training sample used to construct $\widehat{g}$.
The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{0})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes P_{\mathbf{X},\mathbf{Z}})}$. Consequently, the estimator converges in $L^{2}$ over the entire time domain $(0,\tau)$ with respect to Lebesgue measure, and not merely at observed event times or in a way that depends on the censoring pattern.
The next result shows that the rate obtained in Theorem 1 is essentially optimal, up to logarithmic factors. It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{P_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
Theorem 2.
Under A1A2 and A5 there exists $c>0$ such that
$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
Next, we turn to the parametric component $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model. To characterize the nuisance tangent space associated with $g_{o}$, we consider differentiable one-dimensional submodels that pass through $g_{o}$. Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ satisfying
$\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|_{L^{2}([0,\tau]\times[0,1]^{r})}=0\,.$
Here, the scalar $b$ is a perturbation parameter indexing the submodel. The set of all attainable first–order directions at $g_{o}$ is then defined as
$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t. }\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|=0\Big\}\,.$
Finally, semiparametric efficiency theory requires a closed linear subspace, so we take the closure of the linear span of $\mathcal{T}{g_{o}}$, denoted $\overline{\mathcal{T}}{g_{o}}$, which forms the nuisance tangent space.
[…]
$\mathbb{E}\!\left\{\Delta\,\|{\bf Z}-{\bf g}^{*}(T,{\bf X})\|_{c}^{2}\right\},\quad\|v\|_{c}^{2}=(v_{1}^{2},\ldots,v_{p}^{2})^{\top}\,$,
so ${{\bf g}}^{*}(t,{\bf X})$ projects the covariates $\bf Z$} onto the tangent space of the nuisance function. Finally, under the counting-process representation of the model (andersen1982cox), the counting process $N(t)=\Delta I(T\leq t)$ admits the Doob-Meyer decomposition where the associated martingale is
$M(t)=N(t)-\int_{0}^{t}Y(s)\exp\{\mbox{${\theta}$}_{o}^{\sf T}{{\bf Z}}+g_{o}(s,{{\bf X}})\}\,ds\,.$
Theorem 3.
Under Assumptions A1A2 and A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (1) is
$\ell_{\theta_{o}}^{*}(T,{\bf X},{\bf Z},\Delta)\;=\;\int_{0}^{\tau}\big\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\big\}\,dM(t)\,$,
where ${\bf g}^{*}(t,{\bf X})$ is the projection of $\bf Z$} onto the nuisance tangent space and $M(t)$ is the associated counting–process martingale. The semiparametric information bound for $\mbox{${\theta}$}_{0}$ is
$I(\mbox{${\theta}$}_{0})\;=\;\mathbb{E}\!\left\{\,\ell_{\theta_{o}}^{*}(T,{\bf X},{\bf Z},\Delta)^{\otimes 2}\right\}=\mathbb{E}\!\left[\,\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right]\,.$
Consequently, the asymptotic variance of any regular asymptotically linear (RAL) estimator of $\mbox{${\theta}$}_{o}$ is bounded below by $I(\mbox{${\theta}$}_{o})^{-1}$ in the Loewner sense.
With the curvature and metric properties of the population loss, the convergence rate from Theorem 1, and the efficient score representation of Theorem 3, we can now establish an asymptotic linear expansion for $\widehat{\mbox{${\theta}$}}$. This, in turn, yields the asymptotic distribution of the estimator.
Theorem 4.
Under Assumptions A1A5, suppose that the efficient information matrix $I(\mbox{${\theta}$}_{o})$ is nonsingular and that $\sqrt{n}\,\tau_{n}^{2}\;\to\;0$ as $n\to\infty$, where $\tau_{n}=\gamma_{n}\log^{2}n$. Then,
$\sqrt{n}(\widehat{\mbox{${\theta}$}}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\,\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})+o_{p}(1)\,$,
and hence
$\sqrt{n}(\widehat{\mbox{${\theta}$}}-\mbox{${\theta}$}_{o})\;\overset{D}{\longrightarrow}\;N({\bf 0},I(\mbox{${\theta}$}_{o})^{-1})\,.$
The key technical observation is that, in the Taylor expansion of the population score around $g_{o}$, the first–order derivative in $\widehat{g}-g_{o}$ direction vanishes because the efficient score is orthogonal to the nuisance tangent space. The remaining second–order term is of smaller order, specifically $o_{p}(n^{-1/2})$, provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
Finally, we justify the numerical approximation used to compute the estimator in practice. The next theorem shows that, under mild conditions the estimator $\widetilde{}\mbox{${\theta}$}$ has the same asymptotic distribution as the oracle estimator $\widehat{}\mbox{${\theta}$}$.
Theorem 5.
[…]
$\delta_{n}\;=\;\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$
denote the mesh size of this grid. Assume that $\delta_{n}=o(\tau_{n}^{2})$ and $\sqrt{n}\,\delta_{n}\;\rightarrow\;0$, as $n\rightarrow\infty$, and that the conditions of Theorems 1 and 4 hold. Then,
  1. 1.
    The numerically integrated likelihood estimator attains the same convergence rate as the oracle estimator:
    $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})\,,\qquad\|\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})\,.$
  2. 2.
    The parametric estimators are asymptotically equivalent:
    $\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\;\xrightarrow{P}\;{\bf 0}\,.$
    Consequently,
[…]
    where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix from Theorem 3.
Theorem 5 shows that the difference between the exact and the numerically integrated likelihood estimators is $o_{p}(n^{-1/2})$, so that both share the same asymptotic normal distribution. Part 2 verifies that the required error bounds hold for a simple equally spaced grid with $n$ time points, where the per–observation quadrature error is of order $1/n$. This ensures that the asymptotic efficiency result for $\widehat{}\mbox{${\theta}$}$ carries over directly to the implementable estimator $\widetilde{}\mbox{${\theta}$}$ based on numerical integration.
5 Pratical Implementation
5.1 Applying DNN
In practice, we approximate $g_{o}$ by a fully connected ReLU network and model the linear term $\mbox{${\theta}$}^{\top}{\bf Z}$ using a parallel linear layer. The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3). This architecture jointly learns the nuisance function $g$ and a finite–dimensional linear effect of $\bf Z$} in a way that mirrors the model structure.
To evaluate the approximated log-likelihood (3) in practice, we discretize the time axis on a fixed grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$. The dataset is then expanded in the usual counting‐process manner. Each subject $i$ with data $(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})$ is represented by one dummy-observation for every interval $t_{k},t_{k+1})$ such that $t_{k}\leq T_{i}$. All dummy-observations inherit $(\mathbf{X}_{i},\mathbf{Z}_{i})$; the event indicator $\Delta_{ik}=\mathbf{1}\{T_{i}\in(t_{k-1},t_{k}],\Delta=1\}$ is set to 0 except for the final interval containing $T_{i}$, where it takes the original value $\Delta_{i}$. This choice of grid and data expansion enabling efficient computation within standard deep-learning frameworks. The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (therneau2000cox; hu2023conditional).
In theory we allow for sparsity, but in practice we do not explicitly enforce it. Network hyperparameters (number of layers, hidden units, activation functions, learning rate, batch size) are selected via tuning on the training set. Training maximizes the empirical likelihood using the ADAM optimizer in TensorFlow. The data are split into training and validation sets, with 33% reserved for early stopping. Training continues until the validation loss stops improving, and the weights achieving the best validation performance are retained. The linear $\mathbf{Z}$-branch is initialized using the coefficients from a standard Cox model fit on $(\mathbf{X},\mathbf{Z})$, providing a good starting point for the linear component.
After training, we extract the weights from the $\mathbf{Z}$-branch as an initial estimate of $\theta$}. To mitigate the effect of early stopping on the optimization of $\theta$}, we then perform a separate numerical maximization of the likelihood with respect to $\boldsymbol{\theta}$, holding $\widetilde{g}$ fixed at its network estimate. Thus, we treat $\widetilde{g}$ as known and solve a low-dimensional optimization problem for $\boldsymbol{\theta}$ using SciPy’s optimization routines. This post-hoc refinement consistently improved the finite-sample behavior of $\widetilde{\mbox{${\theta}$}}$ in our experiments. For the theoretical results, the critical requirement is that $\widehat{g}$ achieves the convergence rate needed for Theorem 4.
5.2 Estimation of the Covariance Matrix $I(\mbox{${\theta}$}_{o})^{-1}$
Inference on $\mbox{${\theta}$}_{0}$ relies an estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$ from Theorem 4, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$. Recall that the efficient information is
$I(\mbox{${\theta}$}_{o})=\mathbb{E}\!\Big[\Delta\{{\bf Z}-{{\bf g}}^{\ast}(T,{\bf X})\}\{{\bf Z}-{\bf g}^{\ast}(T,{\bf X})\}^{\top}\Big]\,$,
where $\bf g$$^{\ast}$ is the event–weighted projection of $\bf Z$} onto the nuisance tangent space. In practice we estimate $\bf g$$^{\ast}$ using an additional set of neural networks and then construct a plug–in estimate of $I(\mbox{${\theta}$}_{o})$ from the corresponding residuals.
Specifically, we approximate $\bf g$$^{\ast}$ coordinatewise. For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by
$\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}\,.$
Each network $j$, $j=1,\ldots,p$, takes $(T_{i},\mathbf{X}_{i})$ as input, $Z_{ij}$ as target, and uses the same architecture and hyperparameters as the main model; empirically, these choices produce stable estimates and are not sensitive to tuning.
To avoid overfitting, we employ 55–fold cross–fitting. The data are split into five folds, for each fold and each coordinate $j$, the corresponding network is trained on the remaining four folds (restricting to $\Delta=1$) and predictions $\widehat{g}^{\ast}_{j}(T_{i},X_{i})$ are obtained for the held-out fold. for all $i$ in the held–out fold. Aggregating across folds yields, for each observation $i$, the estimated projection vector
$\widehat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})=\bigl(\widehat{g}^{\ast}_{1}(T_{i},{{\bf X}}_{i}),\ldots,\widehat{g}^{\ast}_{p}(T_{i},{{\bf X}}_{i})\bigr)^{\top}\,.$
We then compute the residuals ${\bf R}_{i}={{\bf Z}}_{i}-\hat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into an $n\times p$ matrix $\mathcal{R}$. Our estimator of the efficient information is $\widehat{I}=\mathcal{R}^{\top}\mathcal{R}/n$, and the asymptotic covariance is $\widehat{I}^{-1}$. In practice this procedure is numerically stable, requires no tuning beyond the main network fit, and produces reliable covariance estimates for $\widetilde{\mbox{${\theta}$}}$.
6 Simulation Study
In this section, we compare the finite-sample performance of the proposed FLEXI-Haz method with the partially linear neural PH estimator of zhong2022deep. Our evaluation focuses on the estimation accuracy of the finite-dimensional parameter $\theta$} and on the coverage properties of the corresponding Wald confidence intervals.
We consider a setting with both nuisance and primary covariates. Let ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where each covariate is independently sampled from a uniform distributed over [1,1][-1,1]. The true parameter of interest is $\mbox{${\theta}$}_{o}=(b_{1},b_{2})^{\top}=(2,-1)^{\top}$. For the nuisance component, we define $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects and interactions among the nuisance covariates. The event time $U$ is generated from a hazard function of the form
$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau]\,.$
with $\tau=30$ chosen as the fixed study horizon. Right censoring is introduced through an independent censoring time $C$ sampled from an exponential distribution with expectation $1/30$, combined with administration censoring at time 30. The parameters of the data-generating mechanism are chosen so that the censoring rate is approximately 30%.
For both FLEXI-Haz and the estimator of zhong2022deep, the nuisance function is estimated using a fully connected feedforward network with ReLU activations. The only and difference between the two estimators lies in the underlying hazard model, and the training pipeline and tuning strategy are identical. Hyperparameters are selected once, using only one simulated dataset. We construct the expanded dataset using a fixed time grid of $m_{n}=n$ points augmented with all observed times, and perform 5-fold cross-validation over the following search space:
  • number of hidden layers: $\{1,2,3,4,5,6\}$
  • width (nodes per hidden layer): $\{20,40,80,160,320\}$
  • learning rate: $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$
Training uses early stopping with a patience of 35 epochs and a batch size of $10^{5}$ on the expanded dataset. For each hyperparameter combination we train the network once and record the average validation loss across the five folds; the configuration achieving the smallest average loss is selected. Cross-validation selects, for FLEXI-Haz, a network with five hidden layers, width 20, and learning rate $10^{-3}$. For the estimator of zhong2022deep, the selected architecture has four hidden layers, width 80, and the same learning rate $10^{-3}$.
Table 1 presents the results for $n=8000$ based on 200 replications. The proposed estimator, FLEXI-Haz, is essentially unbiased and its empirical variability closely matches the theoretical standard errors, yielding coverage probabilities near their nominal levels. In contrast, the partially linear PH estimator of zhong2022deep exhibits substantial bias in both components of $\boldsymbol{\theta}_{o}$, leading to near-zero empirical coverage.
Table 1: Simulation results for $n=8000$ observations, based on 200 replications. The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$. “Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90% , 95%” the empirical coverage rates of the corresponding Wald-type confidence intervals.
FLEXI-Haz zhong2022deep
Est Emp SD Est SE 90% 95% Est Emp SD Est SE 90% 95%
$\theta_{o1}$ 1.997 0.040 0.035 0.878 0.935 1.764 0.039 0.034 0.000 0.000
$\theta_{o2}$ -0.994 0.027 0.029 0.937 0.959 -0.886 0.030 0.028 0.015 0.035
7 Discussion
In this work we proposed FLEXI–Haz, a deep semiparametric survival model that combines a fully non-PH formulation with a partially linear structure. By modeling the log-hazard as $\boldsymbol{\theta}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})$ and estimating $g_{o}$ with a deep ReLU network, we retain interpretability and valid inference for the low-dimensional parameter $\boldsymbol{\theta}_{o}$ while allowing highly flexible, time-varying effects of the nuisance covariates. Our main theoretical result is that FLEXI–Haz enjoys the same key properties that zhong2022deep established for partially linear Cox models under PH: the neural estimator of $g_{o}$ attains the optimal composite minimax rate, and the estimator of $\boldsymbol{\theta}_{o}$ is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Crucially, these properties continue to hold in our more general, baseline-free non-PH setting, despite the presence of time in the nuisance component and the use of the full likelihood rather than a partial likelihood.
The simulation study supports these theoretical findings and illustrates their practical relevance. In a setting with nonlinear time–covariate interactions, FLEXI–Haz produces essentially unbiased estimates of $\boldsymbol{\theta}_{o}$, with empirical standard deviations that align closely with the estimated standard errors and nominal coverage for Wald-type confidence intervals. In contrast, the partially linear PH estimator of zhong2022deep, which forces the nuisance component to depend only on $\mathbf{X}$, exhibits substantial bias and severe undercoverage when the true hazard is non-proportional. This behavior is exactly what the theory predicts: when time-varying effects are incorrectly constrained to be time-invariant, the nuisance approximation error leaks into the finite-dimensional parameter. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ removes this source of misspecification, and the empirical results confirm that the resulting estimator behaves in accordance with the asymptotic theory.
Although FLEXI–Haz provides consistent estimators of the hazard and enables estimation of the conditional survival function $S(t\mid\mathbf{X},\mathbf{Z})$, the present work does not derive the asymptotic distribution of the resulting survival-curve estimator. Nevertheless, valid pointwise confidence intervals and simultaneous confidence bands can still be constructed by applying the ensemble–bootstrap procedure developed by arie2024confidence. That method is fully nonparametric, accommodates right censoring, and can be used as a post–hoc procedure on top of FLEXI–Haz without modifying the training algorithm.
Our implementation focuses on fully connected ReLU networks, which fit naturally into the composite Hölder framework and for which sharp approximation and entropy bounds are available. This choice enables us to obtain clean minimax rates and empirical-process control. At the same time, fully connected architectures are not always the most natural representation for complex structured data such as images. Convolutional neural networks (CNNs) are often preferable in imaging applications, and several deep survival methods in the literature rely on convolutional encoders. While the FLEXI–Haz likelihood and estimation strategy can be used with convolutional backbones in practice, our current theory does not cover convolutional architectures, and the existing results on sparse ReLU networks cannot be directly transferred to CNNs. Extending the approximation and efficiency analysis to convolutional or other structured architectures is therefore an important direction for future work.
A further practical advantage of our formulation is that it meshes naturally with the standard “pseudo-observation” representation used to handle time-dependent covariates. The full-likelihood derivation already requires us to expand each subject into multiple pseudo-observations over risk intervals. This is precisely the data-expansion scheme commonly used to encode time-dependent covariates in Cox-type models. As a result, extending FLEXI–Haz to time-dependent covariates is conceptually straightforward: one augments $\mathbf{X}$ to include time-varying components and constructs the expanded dataset in the usual way. The theoretical development is also compatible with this extension, as the key arguments are expressed in terms of predictable processes and integrals over the at-risk time, and do not rely on time-independence of the covariates beyond notational convenience.
The main current limitation of FLEXI–Haz is computational rather than statistical. Because we work with the full likelihood and approximate the cumulative hazard via a Riemann-type sum on a fine time grid, the expanded dataset contains $O(n\,m_{n})$ pseudo-observations, where $m_{n}$ is the number of grid points. In the default choice $m_{n}\asymp n$, this leads to an effective $O(n^{2})$ cost per epoch, which can become prohibitive for very large studies or dense event processes. Our experiments show that this is manageable for moderate $n$, but improving scalability is crucial for applications with hundreds of thousands of subjects or very fine time resolution.
Several strategies could mitigate this computational burden. One natural approach is to sub-sample pseudo-observations corresponding to non-events in the risk sets, e.g., by keeping all intervals containing failures and a carefully chosen random subset of non-event intervals, with importance weights to preserve unbiasedness. Since the information about $\boldsymbol{\theta}_{o}$ and $g_{o}$ is concentrated near failures and in regions with high hazard, aggressive sub-sampling of “uninformative” background intervals should reduce cost with limited loss in efficiency (Keret02102023). Another line of work is to exploit structured quadrature rules or other powerful approximation methods.
Our theoretical treatment of numerical integration in Theorem 5 is deliberately abstract and suggests a broad class of admissible approximations. The only requirement is that the approximation error in the empirical loss is $O_{p}(n^{-1})$ uniformly over the parameter space, and that the corresponding error in the $\boldsymbol{\theta}$-score is $O_{p}(n^{-1/2})$. Any numerical scheme—whether based on fewer grid points, higher-order quadrature, adaptive time meshes, or stochastic approximations—that satisfies these two error bounds yields the same asymptotic distribution of $\widetilde{\mbox{${\theta}$}}$. In particular, this shows that our choice of a simple Riemann sum with $m_{n}\asymp n$ is sufficient but not necessary: one can potentially use substantially fewer time points while preserving the validity of the asymptotic theory, provided the approximation quality is high enough.
8 Appendix - Proofs
See 1
Proof of Lemma 1.
Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (1). Because $f(t,{\bf X},{\bf Z})$ is predictable, the stochastic integral $\int_{0}^{\tau}f(t,{{\bf X}},{{\bf Z}})\,dM(t)$ has expectation zero. Consequently,
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
Since $\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)=\Delta\,f(T,{\bf X},{\bf Z})$, we obtain the identity
$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (10)
Under Assumptions A1–A2, the hazard is uniformly bounded and bounded away from zero,
$0<c_{\lambda}\;\leq\;h_{0}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{\lambda}<\infty.$
Applying these bounds to the right-hand side of (10) yields
$c_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
This implies the desired equivalence:
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\asymp\mathbb{E}\!\{\Delta\,f(T,{\bf X},{\bf Z})\}\,.$
See 2
Proof of Lemma 2.
By the tower property of conditional expectation,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}\Big\{\phi(T,{\bf X},{\bf Z})^{2}\,\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\Big\}\,.$
Assumption A4 states that
$0<\delta_{1}\;\leq\;\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\;\leq\;1\qquad\text{a.s}.$
Since $\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, we obtain
$\delta_{1}\,\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\,\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Rearranging the inequalities gives
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\frac{1}{\delta_{1}}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Thus (7) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
Next, we relate the event–weighted quantity $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}$ to the time–integrated $L^{2}$ norm of $\phi$. Write,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (11)
and let $f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})$ denote the conditional density of $T$ given $({\bf X},{\bf Z})$ on the event $\{\Delta=1\}$. Then,
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}=\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\,dt\,.$
Under Assumptions A1 and A2, the event-time hazard $h_{0}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity. Similarly, Assumption A4 implies the censoring mechanism is well behaved. Therefore,
$\Pr(T>t\mid{\bf X},{\bf Z})=\exp\left[-\int_{0}^{t}\{h_{o}(s\mid{\bf X},{\bf Z})+\mu(s\mid{\bf X},{\bf Z})ds\}\right]$
where $\mu(t\mid{\bf X}.{\bf Z})$ is the censoring hazard bounded away from zero and infinity, and
$f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})=h_{o}(s\mid{\bf X},{\bf Z})\exp\left[-\int_{0}^{t}\{h_{o}(s\mid{\bf X},{\bf Z})+\mu(s\mid{\bf X},{\bf Z})ds\}\right]$
is bounded above and below by positive constants that depend only on $(\tau,\delta_{1},\delta_{2},c_{h},C_{h})$. Specifically, there exist $0<c_{f}\leq C_{f}<\infty$ such that, almost surely,
$c_{f}\;\leq\;f_{T,\Delta=1\mid{\bf X},{\bf Z}}(t\mid{\bf X},{\bf Z})\;\leq\;C_{f},\qquad t\in[0,\tau]\,.$
Using these bounds inside the conditional expectation (11), we obtain, that given $({\bf X},{\bf Z})$,
$c_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\;\leq\;C_{f}\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\,.$
Taking expectation over $({\bf X},{\bf Z})$ yields
$c_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;C_{f}\,\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}.$
Thus (8) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
Finally, suppose $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$. Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. on $\{\Delta=1\}$. From (7) we also have
$0\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–almost surely. Likewise, from (8),
$0\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=0\,$,
so
$\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}=0\,.$
Since the integrand $\phi(t,{\bf X},{\bf Z})^{2}$ is nonnegative, this implies
$\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt=0\quad\text{for $P_{{\bf X},{\bf Z}}$–a.e. }({\bf X},{\bf Z}),$
and hence $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue–a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$. Equivalently,
$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes P_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
See 3
Proof of Lemma 3.
Recall that $\ell_{0}(\eta)$ denotes the population (expected) log-likelihood,
$\ell_{0}(\eta)=\mathbb{E}\Big\{\,\Delta\,\chi_{\eta}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta}(t,{\bf X},{\bf Z})\,dt\Big\},$
where $\chi_{\eta}$ is the linear predictor and $h_{\eta}(t,{\bf X},{\bf Z})$ is the corresponding instantaneous hazard process. Let $\eta_{0}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{0}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
$\eta_{\varepsilon}=\eta_{0}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$. Because $\chi_{\eta}$ is linear in $(g,\mbox{${\theta}$})$, the Gateaux derivative of $\chi_{\eta}$ in direction $w$ at $\eta_{0}$ is simply
$\dot{\chi}_{h}(t,{\bf X},{\bf Z})=\frac{d}{d\varepsilon}\,\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,$,
and we can write
$\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=\chi_{\eta_{0}}(t,{\bf X},{\bf Z})+\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,.$
For notational simplicity, when we fix a particular $\eta$, we also write
$q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})\,$,
so $q$ is the difference in linear predictors between $\eta$ and $\eta_{o}$. Along the path $\eta_{\varepsilon}$, $q$ is proportional to the direction $\dot{\chi}_{w}$.
Next, note that
$h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})\}=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
so at $\varepsilon=0$ we have $h_{\eta_{0}}=h_{o}$ and
$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
By the definition of the Gateaux derivative,
$\ell_{0}^{\prime}(\eta_{o})[w]=\frac{d}{d\varepsilon}\,\ell_{0}(\eta_{\varepsilon})\Big|_{\varepsilon=0}\,.$
Using the explicit form of $\ell_{0}$ and differentiating under the expectation,
$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
Since
$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dN(t)\,$,
the Doob–Meyer decomposition for the true model (1) gives
$dN(t)=Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+dM(t)\,$,
where $M(t)$ is a martingale. Therefore,
$\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})$ $=\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)\,.$
Substituting this into the expression for $\ell_{0}^{\prime}(\eta_{o})[w]$ and cancelling the compensator term,
$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\mathbb{E}\Big\{\int_{0}^{\tau}\dot{\chi}_{h}(t,{\bf X},{\bf Z})\,dM(t)\Big\}\,.$
The stochastic integral $\int_{0}^{\tau}\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dM(t)$ is a martingale with mean zero. Hence,
$\ell_{0}^{\prime}(\eta_{0})[w]=0\qquad\text{for every direction }w\,.$
In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$, we can summarize this step as
$\ell_{0}^{\prime}(\eta_{0})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t,{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
To control $\ell_{0}(\eta)-\ell_{0}(\eta_{0})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$. Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{0}+v(\eta-\eta_{0})$ for $v\in[0,1]$. The mean value theorem yields
$\ell_{0}(\eta)=\ell_{0}(\eta_{0})+\ell_{0}^{\prime}(\eta_{0})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{0}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{0})[w]=0$, it remains to analyse the second derivative term.
For a fixed $v\in[0,1]$, consider the path $\varepsilon\mapsto\eta_{v}+\varepsilon w$ and write
$q_{v}(t,{\bf X},{\bf Z})=\frac{d}{d\varepsilon}\,\chi_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,$,
which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{0}$ gives
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ $=\frac{d^{2}}{d\varepsilon^{2}}\,\ell_{0}(\eta_{v}+\varepsilon w)\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\frac{d^{2}}{d\varepsilon^{2}}\chi_{\eta_{v}+\varepsilon h}(T,{\bf X},{\bf Z})\Big|_{\varepsilon=0}-\int_{0}^{\tau}Y(t)\,\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}\,.$
Since $\chi_{\eta}$ is linear in $\eta$, its second derivative in $\varepsilon$ vanishes, while the second derivative of the hazard is
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t,{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
Hence
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
For notational simplicity (since we will ultimately evaluate this at $\varepsilon=1$), we drop the subscript $v$ and simply write $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$ along the line, so that
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
Next, for relating the second derivative to $d(\eta,\eta_{0})^{2}$, we apply Lemma 1 for
$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\,.$
Indeed, Lemma 1 gives
$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big\}\,.$
Under Assumptions A1A2, the processes $\chi_{\eta}$ are uniformly bounded, and hence
$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
Therefore,
$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Combining this with the expression for $\ell_{0}^{\prime\prime}(\eta_{v})[w,w]$ yields the two-sided bound
$C_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}.$
Recall that the distance $d(\eta,\eta_{0})$ is defined by
$d(\eta,\eta_{0})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{0}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
Hence
$C_{\star}\,d(\eta,\eta_{0})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{0})^{2}\,.$
Therefore, by the Taylor expansion with the mean value theorem,
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{0})^{2}$. Thus,
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})\;\asymp\;-\,d(\eta,\eta_{0})^{2},$
which completes the proof. ∎
In the following, $\mathbb{P}_{n}$ and $\mathbb{P}$ denote the empirical and probability measures of $(T,\Delta,{\bf X},{\bf Z})$, and $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$ is the empirical process. The symbol $\mathbb{E}^{\ast}$ stands for outer expectation, used to avoid measurability issues in the supremum over an uncountable class of functions.
Lemma 4.
Consider the following classes of functions indexed by the parameter $\eta$:
  • The log-likelihood increment
    $f_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\Delta\,\chi_{\eta}(T,\mathbf{X},\mathbf{Z})-\int_{0}^{T}Y(t)\,\exp\{\chi_{\eta}(t,\mathbf{X},\mathbf{Z})\}\,dt,$
  • The efficient score-type term
    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-g^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
    where $g^{\ast}$ denotes the efficient projection defined in Theorem 3.
For $0<\delta\leq 1$, define the localized classes
$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,.$
Under Assumptions A1A3 and A5, there exists a constant $C>0$ such that for every $\delta\in(0,1]$,
$\mathbb{E}^{\ast}\Bigl\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}\bigl|\sqrt{n}\,(\mathbb{P}_{n}-\mathbb{P})\,\mu\bigr|\Bigr\}\;\leq\;C\Bigl\{\delta\,\sqrt{s\log(\mathcal{U}/\delta)}+n^{-1/2}\,s\log(\mathcal{U}/\delta)\Bigr\},$
where $s$ is the sparsity of the neural network class, and
$\mathcal{U}=K\prod_{k=0}^{K}(p_{k}+1)\sum_{k=0}^{K}p_{k}p_{k+1}.$
Proof of Lemma 4.
Fix $\eta_{1},\eta_{2}$ and denote $q(t,{\bf X},{\bf Z})=\chi_{\eta_{1}}(t,{\bf X},{\bf Z})-\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$. For notational convenience, we write $\chi_{\eta}(t)=\chi_{\eta}(t,{\bf X},{\bf Z})$ and $q(t)=\chi_{\eta_{1}}(t)-\chi_{\eta_{2}}(t)$. Because $\chi_{\eta}$ is uniformly bounded under A1–A2 and depends on the network parameters in a Lipschitz way, we have
$\bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\bigr|\;\lesssim\;|q(t,{\bf X},{\bf Z})|\,.$
Using the definition of $f_{\eta}$ and the triangle and Cauchy–Schwarz inequalities,
$|f_{\eta_{1}}-f_{\eta_{2}}|\leq\Delta\,|q(T)|+\int_{0}^{\tau}Y(t)\,|\exp\{\chi_{\eta_{1}}(t)\}-\exp\{\chi_{\eta_{2}}(t)\}|\,dt\,.$
Plugging the bound, squaring and applying the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ gives
$(f_{\eta_{1}}-f_{\eta_{2}})^{2}\;\lesssim\;2\Delta\,q(T)^{2}+2\int_{0}^{\tau}Y(t)\,q(t)^{2}\,dt,$
where all quantities are uniform in $({\bf X},{\bf Z})$ by A1A2. By Lemma 1 we get
$E\!\left[\int_{0}^{\tau}Y(t)\,q(t)^{2}\,dt\right]\asymp E\bigl[\Delta\,q(T)^{2}\bigr]\,.$
The right-hand side is exactly the squared metric $d^{2}(\eta_{1},\eta_{2})$. Thus,
$\|f_{\eta_{1}}-f_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Repeating the argument for $\psi_{\eta}$ requires only replacing the exponential term with the stochastic integral. Assumption A5 ensures boundedness of the predictable integrand ${\bf Z}-g^{\ast}(t,{\bf X})$ and square-integrability of the martingale increments; A1A2 control the compensator. The same calculation shows
$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$. The parametrization $\eta=(\mbox{${\theta}$},g)$ is $s$-sparse in the neural-network coordinates, and Assumption A4 ensures that the weights are uniformly bounded and that the architecture has fixed depth and layer widths. In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in zhong2022deep. Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(P)$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5. Since Lipschitz images of a function class cannot increase its covering number by more than a multiplicative constant, the metric entropy of the localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ is therefore controlled by the entropy of the underlying sparse neural-network ball
$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\}.$
Lemma 6 of zhong2022deep establishes that for such a sparse network class,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{H}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right),$
where $\mathcal{U}$ is the architectural complexity factor. Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union. Consequently,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
Next, we bound the bracketing entropy integral. Using the entropy estimate derived above, the associated Dudley integral satisfies
$J_{[]}(\delta)=\int_{0}^{\delta}\!\sqrt{\log N_{[]}(\varepsilon)}\;d\varepsilon\;\leq\;C_{2}\,\delta\,\sqrt{s\log\!\left(\frac{\mathcal{U}}{\delta}\right)}\,.$
Finally, applying Lemma 3.4.2 of vdvwellner1996 to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(P)}\,\log N_{[]}(\delta)\Bigr\}\,.$
Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(P)}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality. ∎
The following lemma establishes equivalence of expectations under two models whose hazard functions differ only through the nuisance function $g$. Because the hazards are uniformly bounded above and below, the distributions under $g_{0}$ and $g_{1}$ are mutually absolutely continuous with Radon–Nikodym derivatives uniformly bounded. This equivalence allows us to control expectations under one model using expectations under another.
Lemma 5.
Suppose Assumptions A1A2 and A4 hold. Let $g_{0},g_{1}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and fix $\mbox{${\theta}$}\in\Theta$. For $j=0,1$, let $P_{g_{j}}$ denote the distribution of $(T,{\bf X},{\bf Z},\Delta)$ under event hazard $h_{g_{j}}(t\mid{\bf X},{\bf Z})=\exp\{g_{j}(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}$ and a common censoring hazard $\mu(t\mid{\bf X},{\bf Z})$. As a consequence of Assumptions A1A2, there exists a constant $0<D<\infty$ such that, almost surely,
$e^{-D}\;\leq\;h_{g_{j}}(t\mid\boldsymbol{X},\boldsymbol{Z})\;\leq\;e^{D}\,,\qquad\mbox{for all}\,\,t\in[0,\tau]\,,\ j=0,1\,.$
Define $c=e^{-D}$ and $C=e^{D}$. Then, for any nonnegative measurable function $q:[0,\tau]\times\mathcal{X}\to[0,\infty)$ with $\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]<\infty$,
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\!\big\{q(T,{\bf X})\big\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\!\big\{q(T,{\bf X})\big\}\,$,
where
$K_{-}=\frac{c}{C}\,e^{-\tau(C-c)}\,,\qquad K_{+}=\frac{C}{c}\,e^{\,\tau(C-c)}\,.$
In particular, $\mathbb{E}_{P_{g_{1}}}[q(T,{\bf X})]\asymp\mathbb{E}_{P_{g_{0}}}[q(T,{\bf X})]$ with constants depending only on $(c,C,\tau)$.
Proof of Lemma 5.
We work conditional on the covariates $({\bf x},{\bf z})$ and then integrate over their distribution. Under model $g_{j}$, the event time $U_{j}$ has density and survival
$f^{U}_{j}(t\mid{\bf x},{\bf z})=h_{j}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z}),\qquad S^{U}_{j}(t\mid{\bf x},{\bf z})=\exp\!\Big\{-\!\int_{0}^{t}h_{j}(s\mid{\bf x},{\bf z})\,ds\Big\}.$
The censoring time $C$ has density $f^{C}(t\mid{\bf x},{\bf z})=\mu(t\mid{\bf x},{\bf z})S^{C}(t\mid{\bf x},{\bf z})$ and survival $S^{C}(t\mid{\bf x},{\bf z})$, common to both models. The observed time and event indicator are $T_{j}=\min(U_{j},C)$ and $\Delta_{j}=\mathbf{1}\{U_{j}\leq C\}$. Given $({\bf x},{\bf z})$, their joint density under $g_{j}$ is
$f_{j}(t,\delta\mid{\bf x},{\bf z})=\begin{cases}h_{j}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z})\,S^{C}(t\mid{\bf x},{\bf z})\,,&\delta=1\\[3.99994pt] \mu(t\mid{\bf x},{\bf z})\,S^{C}(t\mid{\bf x},{\bf z})\,S^{U}_{j}(t\mid{\bf x},{\bf z})\,,&\delta=0\,.\end{cases}$
By Assumptions A1A2,
$ct\leq\int_{0}^{t}h_{j}(s\mid{\bf x},{\bf z})\,ds\leq Ct,\qquad e^{-Ct}\leq S^{U}_{j}(t\mid{\bf x},{\bf z})\leq e^{-ct}.$
The likelihood ratio of $(T_{j},\Delta_{j})$ under $g_{1}$ versus $g_{0}$, conditional on $({\bf x},{\bf z})$, is
$L(t,\delta\mid{\bf x},{\bf z})=\frac{f_{1}(t,\delta\mid{\bf x},{\bf z})}{f_{0}(t,\delta\mid{\bf x},{\bf z})}=\left\{\frac{h_{1}(t\mid{\bf x},{\bf z})}{h_{0}(t\mid{\bf x},{\bf z})}\right\}^{\delta}\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}.$
Since $c\leq h_{j}\leq C$,
$\frac{c}{C}\leq\frac{h_{1}(t\mid{\bf x},{\bf z})}{h_{0}(t\mid{\bf x},{\bf z})}\leq\frac{C}{c}.$
Moreover,
$\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}=\exp\!\Big[-\!\int_{0}^{t}\{h_{1}(s\mid{\bf x},{\bf z})-h_{0}(s\mid{\bf x},{\bf z})\}\,ds\Big],$
and because $|h_{1}-h_{0}|\leq C-c$,
$e^{-(C-c)t}\leq\frac{S^{U}_{1}(t\mid{\bf x},{\bf z})}{S^{U}_{0}(t\mid{\bf x},{\bf z})}\leq e^{(C-c)t}.$
For $t\leq\tau$ this yields
$\frac{c}{C}\,e^{-(C-c)\tau}\;\leq\;L(t,\delta\mid{\bf x},{\bf z})\;\leq\;\frac{C}{c}\,e^{(C-c)\tau}.$
Thus the Radon–Nikodým derivative satisfies
$K_{-}\;\leq\;\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\;\leq\;K_{+},\qquad P_{g_{0}}\text{--a.s.},$
where $K_{-}=\frac{c}{C}e^{-\tau(C-c)}$ and $K_{+}=\frac{C}{c}e^{\tau(C-c)}$. For any nonnegative measurable function $q$ with finite expectation under $P_{g_{0}}$,
$\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}=\mathbb{E}_{P_{g_{0}}}\!\left\{q(T,{\bf X})\,\frac{dP_{g_{1}}}{dP_{g_{0}}}(T,\Delta,{\bf X},{\bf Z})\right\}$
and the bounds on the likelihood ratio give
$K_{-}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\;\leq\;\mathbb{E}_{P_{g_{1}}}\{q(T,{\bf X})\}\;\leq\;K_{+}\,\mathbb{E}_{P_{g_{0}}}\{q(T,{\bf X})\}\,.$
Lemma 6.
Assume A2. Let $\eta_{j}=(\mbox{${\theta}$}_{j},g_{j})$, $j=1,2$, be two parameter values in $(\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M))\cup(\Theta\times\mathcal{G}(K,s,\boldsymbol{p},D))$. Define
$\chi_{\eta_{j}}(t,{\bf X},{\bf Z})=g_{j}(t,{\bf X})+\mbox{${\theta}$}_{j}^{\top}{\bf Z},\qquad j=1,2.$
Assume that $\|{\bf Z}\|\leq B_{\bf Z}$ a.s. for some finite constant $B_{\bf Z}$. Then
$\mathbb{E}\Bigl[\Delta\bigl\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\bigr\}^{2}\Bigr]\;\leq\;2\,d(\eta_{1},\eta_{2})^{2}\;+\;2B_{\bf Z}^{2}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Proof of Lemma 6.
Fix $(T,\Delta,{\bf X},{\bf Z})$ and set $a=g_{1}(T,{\bf X})-g_{2}(T,{\bf X})$, $b=(\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2})^{\top}{\bf Z}$. Then $\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})=a+b$. Use the elementary inequality
$(u-v)^{2}\;\leq\;2u^{2}+2v^{2}\qquad\text{for all }u,v\in\mathbb{R},$
with $u=a+b$ and $v=b$. This gives
$a^{2}=(a+b-b)^{2}\;\leq\;2(a+b)^{2}+2b^{2}.$
Multiplying by $\Delta\in\{0,1\}$ preserves the inequality
$\Delta\,a^{2}\;\leq\;2\Delta\,(a+b)^{2}+2\Delta\,b^{2}\,.$
Taking expectations and using the definition of $d(\eta_{1},\eta_{2})$,
$\mathbb{E}\Bigl[\Delta\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\}^{2}\Bigr]$ $=\mathbb{E}(\Delta a^{2})$
$\leq 2\,\mathbb{E}\{\Delta(a+b)^{2}\}+2\,\mathbb{E}(\Delta b^{2})$
$=2\,d(\eta_{1},\eta_{2})^{2}+2\,\mathbb{E}(\Delta b^{2}).$
For the second term, by boundedness of $\bf Z$},
$|b|=|(\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2})^{\top}{\bf Z}|\leq\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|\,\|{\bf Z}\|\leq B_{\bf Z}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|,$
and hence
$b^{2}\;\leq\;B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
Since $\Delta\leq 1$,
$\mathbb{E}(\Delta b^{2})\leq\mathbb{E}(b^{2})\leq B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Putting everything together,
$\mathbb{E}\Bigl[\Delta\{g_{1}(T,{\bf X})-g_{2}(T,{\bf X})\}^{2}\Bigr]\leq 2\,d(\eta_{1},\eta_{2})^{2}+2B_{\bf Z}^{2}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Lemma 7.
If Assumptions A1A5 hold, then there exists a constant $c_{\theta}>0$ depending only on the constants in A1A4 such that, for all parameters in a neighborhood of $\eta_{o}$,
$d(\eta_{1},\eta_{2})^{2}\;\geq\;c_{\theta}\,\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}\,.$
Proof of Lemma 7.
Let $\mbox{${\beta}$}=\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}$, $a(t,{\bf X})=g_{1}(t,{\bf X})-g_{2}(t,{\bf X})$, $b=\mbox{${\beta}$}^{\top}{\bf Z}$. At a failure time, $\chi_{\eta_{1}}-\chi_{\eta_{2}}=a+b$. Thus
$d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\{\Delta(a+b)^{2}\}\,.$
To find the smallest possible value of this expression over all possible nuisance differences $g_{1}-g_{2}$, minimize $(a+b)^{2}$ over scalar functions $a(T,{\bf X})$. The minimizing choice is
$a^{\ast}(T,{\bf X})=-\mathbb{E}_{0}\{b\mid T,{\bf X},\Delta=1\}\,.$
Substituting this optimal choice yields
$\inf_{g_{1}-g_{2}}d(\eta_{1},\eta_{2})^{2}=\mathbb{E}_{0}\!\Big\{\Delta\,Var(b\mid T,{\bf X},\Delta=1)\Big\}\,.$
This is exactly the standard least–squares argument. Assumption A4 ensures
$\Pr(\Delta=1\mid T,{\bf X})\;\geq\;c_{\Delta}\quad\text{a.s.}$
Therefore
$\mathbb{E}_{0}\Big[\Pr(\Delta=1\mid T,{\bf X})\,Var(b\mid T,{\bf X},\Delta=1)\Big]\;\geq\;c_{\Delta}\,\mathbb{E}_{0}\bigl[Var(b\mid T,{\bf X},\Delta=1)\bigr]\,.$
Assumption A5 implies that for some $c_{Z}>0$,
$Var({\bf Z}\mid T,{\bf X},\Delta=1)\succeq c_{Z}I_{p}\,$,
thus for any $\beta$},
$Var(b\mid T,{\bf X},\Delta=1)=Var(\mbox{${\beta}$}^{\top}{\bf Z}\mid T,{\bf X},\Delta=1)\geq c_{\Delta}c_{Z}\|\mbox{${\beta}$}\|^{2}\,.$
Combining the above, we get
$d(\eta_{1},\eta_{2})^{2}\geq c_{Z}\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|^{2}.$
Setting $c_{\theta}=c_{\Delta}c_{Z}>0$ completes the proof. ∎
See 1
Proof of Theorem 1.
Let $\ell(\eta;{\bf W})$, ${\bf W}=(T,\Delta,{\bf X}^{\top},{\bf Z}^{\top})^{\top}$, denote the per–observation log–likelihood contribution, so that $\ell_{n}(\eta)=\mathbb{P}_{n}\ell(\eta;\cdot)$ is the empirical loss and $\ell_{0}(\eta)=\mathbb{P}_{0}\ell(\eta;\cdot)$ is the population loss. Let $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P}_{0})$ be the empirical process. Recall the distance
$d(\eta_{1},\eta_{2})\;=\;\Big(\mathbb{E}_{0}\big[\Delta\{\chi_{\eta_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{2}}(T,{\bf X},{\bf Z})\}^{2}\big]\Big)^{1/2},\qquad\eta=(\mbox{${\theta}$},g),$
and let $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$ denote the true parameter. By Lemma 4, there exists a constant $C>0$ such that for every $0<\delta\leq 1$,
$\mathbb{E}^{*}\!\sup_{d(\eta,\eta_{o})\leq\delta}\bigl||\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|\;\leq\;\Psi_{n}(\delta)\equiv\delta\,\sqrt{s\log\frac{\mathcal{U}}{\delta}}+\frac{s}{\sqrt{n}}\log\frac{\mathcal{U}}{\delta}\,$, (12)
and by Markov’s inequality, this implies
$\sup_{d(\eta,\eta_{o})\leq\delta}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\Psi_{n}(\delta)\bigr)\,.$
Set $\tau_{n}=\gamma_{n}\log^{2}n$ as in the theorem. By Assumption A3 and the specific choice of $(K,s,\boldsymbol{p})$ (cf. the construction in schmidthieber2020nonparametric and zhong2022deep), we have
$s\log\!\Bigl(\frac{\mathcal{U}}{\tau_{n}}\Bigr)=O\!\bigl(n\,\gamma_{n}^{2}\log n\bigr)\,$,
so that
[…]
for $n$ large enough. Hence, there exists a constant $c>0$ such that
$\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\sqrt{n}\,\tau_{n}^{2}\quad\mbox{and}\quad\frac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\;\leq\;\frac{c}{2}\,\tau_{n}^{2}\,.$ (13)
Consequently,
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|=\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\frac{1}{\sqrt{n}}\bigl|\mathbb{G}_{n}\big\{\ell(\eta;\cdot)-\ell(\eta_{o};\cdot)\big\}\bigr|=O_{p}\bigl(\tfrac{1}{\sqrt{n}}\Psi_{n}(\tau_{n})\bigr)=O_{p}(\tau_{n}^{2}),$ (14)
where we used (12) and (13).
By Theorem 1 of schmidthieber2020nonparametric, applied to the composite Hölder class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{\bf d},M)$ with the choice of $(K,s,\boldsymbol{p},D)$ given in Assumption A3, there exists $\tilde{g}_{1}\in\mathcal{G}(K,s,\boldsymbol{p},D)$ such that
$\mathbb{E}\bigl\{\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\leq C_{1}\,\tau_{n}^{2}\,,\qquad\mathbb{E}\bigl\{\Delta\big(\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\big)^{2}\bigr\}\leq C_{1}\,\tau_{n}^{2}$
for some constant $C_{1}>0$ independent of $n$. Define $\tilde{\eta}_{1}=(\mbox{${\theta}$}_{o},\tilde{g}_{1})$. Then,
$d^{2}(\tilde{\eta}_{1},\eta_{o})=\mathbb{E}\bigl[\Delta\big\{\chi_{\tilde{\eta}_{1}}(T,{\bf X},{\bf Z})-\chi_{\eta_{o}}(T,{\bf X},{\bf Z})\big\}^{2}\bigr]=\mathbb{E}\bigl[\Delta\big\{\tilde{g}_{1}(T,{\bf X})-g_{o}(T,{\bf X})\big\}^{2}\bigr]\leq C_{1}\,\tau_{n}^{2}\,$,
since the $\theta$}–component is the same in both $\tilde{\eta}_{1}$ and $\eta_{o}$. Therefore,
$d(\tilde{\eta}_{1},\eta_{o})\;\leq\;\sqrt{C_{1}}\,\tau_{n}\,$, (15)
and in particular $\tilde{\eta}_{1}$ lies inside a ball of radius proportional to $\tau_{n}$ around $\eta_{o}$.
Lemma 3 states that, for any $\eta\in\mathcal{F}_{D}$,
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2}\,$,
that is, there exist constants $0<c_{-},c_{+}<\infty$ such that
$c_{+}\,d(\eta,\eta_{o})^{2}\;\leq\;\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\,.$
Applying this to $\eta=\tilde{\eta}_{1}$ and using (15), we obtain
$\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\;\geq\;-\,c_{+}\,d(\tilde{\eta}_{1},\eta_{o})^{2}\;\geq\;-\,c_{+}C_{1}\,\tau_{n}^{2}\;=\;-\,C_{2}\,\tau_{n}^{2}\,$, (16)
for some constant $C_{2}>0$.
By combining (14) and (16) we get
$\ell_{n}(\tilde{\eta}_{1})-\ell_{n}(\eta_{o})$ $=\bigl\{\ell_{n}(\tilde{\eta}_{1})-\ell_{0}(\tilde{\eta}_{1})\bigr\}-\bigl\{\ell_{n}(\eta_{o})-\ell_{0}(\eta_{o})\bigr\}+\bigl\{\ell_{0}(\tilde{\eta}_{1})-\ell_{0}(\eta_{o})\bigr\}$
$\geq-\,\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|\;-\;C_{2}\,\tau_{n}^{2}$
$=-\,O_{p}(\tau_{n}^{2}).$ (17)
[…]
$\ell_{n}(\widehat{\eta})=\max_{\eta}\ell_{n}(\eta)\;\geq\;\ell_{n}(\tilde{\eta}_{1})\;\geq\;\ell_{n}(\eta_{o})-O_{p}(\tau_{n}^{2})\,.$ (18)
On the other hand, by Lemma 3, the population loss satisfies
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\leq\;-\,c_{-}\,d(\eta,\eta_{o})^{2}\quad\text{for all }\eta\in\mathcal{F}_{D}\,.$
The empirical-process bound (14) and the likelihood chain (18) verify the conditions (i)–(ii) of Theorem 3.4.1 in vdvwellner1996 with radius of order $\tau_{n}$. Therefore,
$d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})\,.$
Finally, by writing
$\chi_{\widehat{\eta}}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})=\{\widehat{g}(t,{\bf X})-g_{o}(t,{\bf X})\}+(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})^{\top}{\bf Z}\,$,
[…]
$\mathbb{E}\Bigl[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\Bigr]\;\leq\;C\Bigl(d(\widehat{\eta},\eta_{o})^{2}+\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\Bigr)\,.$ (19)
Applying Lemma 7 with $\eta_{1}=\widehat{\eta}$ and $\eta_{2}=\eta_{o}$ yields
$\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|^{2}\leq c_{\theta}^{-1}d(\widehat{\eta},\eta_{o})^{2}=O_{p}(\tau_{n}^{2})\,.$
Hence, $\|\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}\|=O_{p}(\tau_{n})$, and the right-hand side of (19) is $O_{p}(\tau_{n}^{2})$, which is exactly the claim of Theorem 1. The expectation is taken with respect to a new observation $(T,\Delta,{\bf X},{\bf Z})\sim P_{0}$, while the $O_{p}(\cdot)$ is with respect to the randomness in the training sample used to construct $\widehat{g}$. ∎
See 2
Proof of Theorem 2.
For $g\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and and $\mbox{${\theta}$}_{o}\in\Theta$, let $P_{g,\theta_{o}}$ denote the joint distribution of $(T,\Delta,{\bf X},{\bf Z})$ induced by the hazard $h_{g,\theta_{o}}(t\mid{\bf X},{\bf Z})=\exp\{g(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}$ together with a censoring hazard that is held fixed across all $g$. By the standard Kullback-Leibler (KL) distance for bounded hazards (Assumptions A1A2), there exists $C>0$ such that for any $g_{0},g_{1}\in\mathcal{\mathcal{}}{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\left[\Delta\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\right]\,.$
Since $\Delta\leq 1$,
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C\,n\,\mathbb{E}_{P_{g_{1},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr].$
Lemma 5 yields the equivalence
$\mbox{KL}(P_{g_{1},\theta_{o}}\,\|\,P_{g_{0},\theta_{o}})\;\leq\;C^{\prime}\,n\,\mathbb{E}_{P_{g_{0},\theta_{o}}}\!\bigl[\{g_{1}(T,{\bf X})-g_{0}(T,{\bf X})\}^{2}\bigr]$
for a constant $C^{\prime}$ depending only on A1A2.
By the packing construction for composite Hölder networks (schmidthieber2020nonparametric), there exist functions $g^{(0)},\dots,g^{(N)}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ and constants $c_{1},c_{2}>0$ such that
$\|g^{(j)}-g^{(k)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}\;\geq\;2c_{1}\,\gamma_{n}\,,\qquad j\neq k\,$, (20)
and
$\frac{1}{N}\sum_{j=1}^{N}\|g^{(j)}-g^{(0)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\;\leq\;\frac{c_{2}\,\log N}{n}\,.$ (21)
Applying the KL bound with $g_{0}=g^{(0)}$ and $g_{1}=g^{(j)}$ and the using (21),
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}\!\big(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}}\big)\;\leq\;C^{\prime}\,c_{2}\,\log N\,.$
Choose $\alpha\in(0,1/8)$ so that $C^{\prime}c_{2}\leq\alpha$. Then,
$\frac{1}{N}\sum_{j=1}^{N}\mbox{KL}(P_{g^{(j)},\theta_{o}}\,\|\,P_{g^{(0)},\theta_{o}})\;\leq\;\alpha\log N\,.$
Applying Tsybakov’s Fano inequality to $\{g^{(0)},\dots,g^{(N)}\}$ with separation (20), we obtain that there exists $c>0$ (depending only on $c_{1}$ and $\alpha$) such that for any estimator $\widehat{g}$,
$\sup_{j=0,\dots,N}\ \mathbb{E}_{P_{g^{(j)},\theta_{o}}}\!\Big(\|\widehat{g}-g^{(j)}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Big)\;\geq\;c\,\gamma_{n}^{2}\,.$
Since the supremum over $g_{o}\in\mathcal{H}$ dominates this finite set,
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\Bigl(\|\widehat{g}-g_{o}\|_{L^{2}(P_{g^{(0)},\theta_{o}})}^{2}\Bigr)\;\geq\;c\,\gamma_{n}^{2}\,.$
By Lemma 5, there exists a constant $C_{\mathrm{eq}}>0$ (depending only on A1A2) such that, uniformly over $g_{o}\in\mathcal{H}$ and all sieve estimators $\widehat{g}$,
$\mathbb{E}_{P_{g^{(0)},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\mathrm{eq}}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big].$
Combining this with the previous display yields
$\inf_{\widehat{g}}\ \sup_{g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c}{C_{\mathrm{eq}}}\,\gamma_{n}^{2}.$
Finally, since the supremum in the theorem is also taken over $\mbox{${\theta}$}_{o}\in\Theta$, and the above bound holds for each fixed $\theta_{o}$, we obtain
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;c_{1}\,\gamma_{n}^{2}$ (22)
for some $c_{1}>0$. By Lemma 2 and Assumptions A1, A2, A4, there exists a constant $C_{\Delta}>0$ such that, for all $g_{o}\in\mathcal{H}$ and all $\widehat{g}$,
$\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\leq\;C_{\Delta}\,\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\,.$
Combining this with (22) gives
$\inf_{\widehat{g}}\ \sup_{\theta_{o}\in\Theta,\ g_{o}\in\mathcal{H}}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\big[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\big]\;\geq\;\frac{c_{1}}{C_{\Delta}}\,\gamma_{n}^{2}\,.$
Renaming the constant $c=c_{1}/C_{\Delta}$ completes the proof. ∎
See 3
Proof of Theorem 3.
Let $\mbox{${\theta}$}\in\mathbb{R}^{p}$ and let $\{g_{s}:s\in(-1,1)\}\subset\mathcal{H}_{g_{o}}$ be a one–dimensional regular submodel of $g$, with
$g_{s}=g_{o}+sg,\qquad g=\left.\frac{\partial g_{s}}{\partial s}\right|_{s=0}\in\mathcal{T}_{g_{o}}\,.$
Consider the semiparametric submodel
$\left\{P_{\theta_{o}+s\theta,\;g_{s}}:s\in(-1,1)\right\}\,.$
For a single observation $(T,\Delta,{\bf X},{\bf Z})$ the log-likelihood is
$\ell(\mbox{${\theta}$},g)=\Delta\{g(T,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}-\int_{0}^{\tau}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt.$
Differentiating $\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})$ at $s=0$ gives
$\left.\frac{d}{ds}\ell(\mbox{${\theta}$}_{o}+s\mbox{${\theta}$},g_{s})\right|_{s=0}=\mbox{${\theta}$}^{\top}\dot{\ell}_{\theta_{o}}+\dot{\ell}_{g_{o},g},$
where
$\dot{\ell}_{\theta_{o}}=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}{\bf Z}\,dM(t),$
and
$\dot{\ell}_{g_{o},g}=\Delta g(T,{\bf X})-\int_{0}^{\tau}g(t,{\bf X})\,Y(t)\exp\{g_{o}(t,{\bf X})+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\}\,dt=\int_{0}^{\tau}g(t,{\bf X})\,dM(t)\,.$
Thus the (naive) score for $\mbox{${\theta}$}_{o}$ is $\dot{\ell}_{\theta_{o}}$. To obtain the efficient score, we project this onto the orthocomplement of the nuisance tangent space
$\mathcal{T}_{g_{o}}^{\mathrm{nuis}}=\left\{\dot{\ell}_{g_{o},g}:g\in\mathcal{T}_{g_{o}}\right\},$
but the projection must be taken onto the closure $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$ in $L^{2}(P_{0})$. Hence the efficient score is
$\ell^{*}_{\theta_{o}}=\dot{\ell}_{\theta_{o}}-\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})\,$,
where $\Pi_{g_{o}}$ denotes the $L^{2}(P_{0})$–projection onto $\overline{\mathcal{T}_{g_{o}}^{\mathrm{nuis}}}$. The projection $\Pi_{g_{o}}(\dot{\ell}_{\theta_{o}})$ has the representation $\dot{\ell}_{g_{o},{\bf g}^{*}}$ for a pp–vector ${\bf g}^{*}=(g_{1}^{*},\dots,g_{p}^{*})^{\top}\in(\overline{\mathcal{T}_{g_{o}}})^{p}$ satisfying the orthogonality condition
$\mathbb{E}\!\Big\{(\dot{\ell}_{\theta_{o}}-\dot{\ell}_{g_{o},{\bf g}^{*}})\,\dot{\ell}_{g_{o},g}\Big\}={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Using Lemma 1 of sasieni1992information, this orthogonality is equivalent to
$\mathbb{E}\!\left[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}\,g(T,{\bf X})\right]={\bf 0}\qquad\text{for all }g\in\mathcal{T}_{g_{o}}\,.$
Thus, componentwise, $g_{j}^{*}$ is the (unique) $\Delta$–weighted $L^{2}(P_{0})$ projection of $Z_{j}$ onto $\overline{\mathcal{T}_{g_{o}}}$. Equivalently, $\bf g$$^{*}$ minimizes
$\mathbb{E}\!\left(\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{c}^{2}\right)\,,\qquad{\bf g}\in(\overline{\mathcal{T}_{g_{o}}})^{p}\,.$
Substituting $\bf g$$^{*}$, the efficient score becomes
$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t).$
By the martingale isometry (again Lemma 1 of sasieni1992information), the efficient information matrix is
$I(\theta_{o})=\mathbb{E}\left[\Delta\,\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\right].$
Finally, under the regularity conditions in Assumptions (A1)–(A4), existence and uniqueness of the weighted projection $\bf g$$^{*}$ in $\overline{\mathcal{T}_{g_{o}}}^{p}$ follow from the approximation theorems of stone1985additive and the general semiparametric projection theory in Appendix A.4 of bickel1993efficient. Therefore, both the efficient score and the efficient information matrix are well-defined and given by the formulas above. ∎
See 4
Proof of Theorem 4.
For any $(\mbox{${\theta}$},g)$, define the per–observation score
$r_{\theta,g}(T,{\bf X},{\bf Z},\Delta)=\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}-\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}Y(t)\exp\{g(t,{\bf X})+\mbox{${\theta}$}^{\top}{\bf Z}\}\,dt\,$,
where $\bf g$$^{*}$ is the pp–dimensional minimizer from Theorem 3, which depends only on $(\mbox{${\theta}$}_{o},g_{o})$. By Theorem 3, at the true parameter $(\mbox{${\theta}$}_{o},g_{o})$ this coincides with the efficient score,
$r_{\theta_{o},g_{o}}(T,{\bf X},{\bf Z},\Delta)=\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta),\qquad\mathbb{P}\,r_{\theta_{o},g_{o}}={\bf 0},$
and the efficient information matrix is
$I(\mbox{${\theta}$}_{o})=\mathbb{P}\bigl\{r_{\theta_{o},g_{o}}^{\otimes 2}\bigr\}=\mathbb{E}\!\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
The empirical efficient score at $(\mbox{${\theta}$},g)$ is $\mathbb{P}_{n}r_{\theta,g}$. By the first–order condition for the (sieve) maximum–likelihood estimator, the estimated parameters satisfy
$\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}\,.$
Write
$\mathbb{P}_{n}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}+(\mathbb{P}_{n}-\mathbb{P})\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}\,.$
By Theorem 1, we have $d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})$, where $\eta=(\mbox{${\theta}$},g)$ and $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$. Using the same entropy bound and maximal inequality as in Lemma 4, one shows that the class
$\mathcal{R}_{n}=\bigl\{r_{\theta,g}:d(\eta,\eta_{o})\leq C\tau_{n}\bigr\}$
is stochastically equicontinuous, and in particular
$\sup_{r\in\mathcal{R}_{n}}|(\mathbb{P}_{n}-\mathbb{P})r|=o_{p}(n^{-1/2})\,.$
[…]
$(\mathbb{P}_{n}-\mathbb{P})\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=o_{p}(n^{-1/2})\,.$
Using $\mathbb{P}_{n}r_{\widehat{\theta},\widehat{g}}={\bf 0}$, this identity can be written as
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,P_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2}).$ (23)
Consider the map
$m(\mbox{${\theta}$},g)=\mathbb{P}\,r_{\theta,g}.$
Since $m(\mbox{${\theta}$}_{o},g_{o})={\bf 0}$ and $r_{\theta,g}$ is smooth in $(\mbox{${\theta}$},g)$ in a neighbourhood of $(\mbox{${\theta}$}_{o},g_{o})$, a first–order expansion gives
$m(\widehat{}\mbox{${\theta}$},\widehat{g})-m(\mbox{${\theta}$}_{o},g_{o})=\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]+R_{n},$
where $\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})$ is the derivative of $m$ with respect to $\theta$} at $(\mbox{${\theta}$}_{o},g_{o})$, $\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\cdot]$ is the Gateaux derivative in the gg–direction, and $R_{n}$ is a second–order remainder.
Differentiating $r_{\theta,g_{o}}$ with respect to $\theta$} and using the representation of the efficient score in Theorem 3 together with the martingale isometry, one obtains
$\dot{m}_{\theta}(\mbox{${\theta}$}_{o},g_{o})=\left.\frac{\partial}{\partial\mbox{${\theta}$}}\right|_{\mbox{${\theta}$}=\mbox{${\theta}$}_{o}}\mathbb{P}\,r_{\theta,g_{o}}=-\,I(\mbox{${\theta}$}_{o})\,.$
Hence the contribution of $\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}$ in the expansion is $I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})$.
Fix a deterministic direction $g$ and consider $g_{s}=g_{o}+sg$. Differentiating under the integral sign and applying the martingale compensation identity yields
$\left.\frac{d}{ds}\right|_{s=0}m(\mbox{${\theta}$}_{o},g_{s})=-\,\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]\,.$
By Theorem 3, $\bf g$$^{*}$ is the $\Delta$–weighted $L^{2}(P)$–projection of $\bf Z$} onto the nuisance tangent space $(\overline{\mathcal{T}}_{g_{o}})^{p}$, so
$\mathbb{E}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}g(T,{\bf X})\bigr]={\bf 0}\quad\text{for all }h\in\overline{\mathcal{T}}_{g_{o}}.$
By Assumptions A1A4, $\widehat{g}-g_{o}\in\overline{\mathcal{T}}_{g_{o}}$, and therefore
$\dot{m}_{g}(\mbox{${\theta}$}_{o},g_{o})[\widehat{g}-g_{o}]=0.$
The map $g\mapsto m(\mbox{${\theta}$}_{o},g)$ is twice Fréchet differentiable, and its second derivative is bounded by a constant multiple of $\mathbb{E}[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}]$. Consequently,
$|R_{n}|\;\lesssim\;\mathbb{E}\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]=O_{p}(\tau_{n}^{2}),$
where Theorem 1 gives $\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}=O_{p}(\tau_{n})$. Our assumption $\sqrt{n}\,\tau_{n}^{2}\to 0$ implies $\tau_{n}^{2}=o(n^{-1/2})$, so $R_{n}=o_{p}(n^{-1/2})$.
[…]
$\mathbb{P}\bigl\{r_{\widehat{\theta},\widehat{g}}-r_{\theta_{o},g_{o}}\bigr\}=-\,I(_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(n^{-1/2})\,.$
Substituting this into (23) yields
$I(\mbox{${\theta}$}_{o})(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=-\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2})\,$,
or, equivalently,
$\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o}=I(\mbox{${\theta}$}_{o})^{-1}\,\mathbb{P}_{n}r_{\theta_{o},g_{o}}+o_{p}(n^{-1/2})\,.$
Recalling that $r_{\theta_{o},g_{o}}=\ell^{*}_{\theta_{o}}$ from Theorem 3, we obtain the stochastic expansion
$\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=I(\mbox{${\theta}$}_{o})^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})+o_{p}(1)\,.$
The efficient score $\ell^{*}_{\theta_{o}}$ has mean zero and covariance $I(\mbox{${\theta}$}_{o})$, again by Theorem 3. By the classical central limit theorem,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ell^{*}_{\theta_{o}}(T_{i},{\bf X}_{i},{\bf Z}_{i},\Delta_{i})\;\overset{D}{\longrightarrow}\;N\bigl({\bf 0},I(\mbox{${\theta}$}_{o})\bigr)\,.$
Slutsky’s theorem then gives
$\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})\;\overset{D}{\longrightarrow}\;N\bigl({\bf 0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,
which completes the proof. ∎
Lemma 8.
Let $\delta_{n}=\max_{j}(t_{j}-t_{j-1})$ be the mesh size of the time grid. Under Assumptions A1A4, there exist constants $C_{1},C_{2}>0$ such that, for all $n$ and all $\eta$ in the neighborhood $\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}$,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)\bigr|\leq C_{1}\,\delta_{n}$ (24)
and
$\sup_{\eta\in\mathcal{N}_{n}}\Bigl\|\nabla_{\theta}\widetilde{\ell}_{n}(\eta)-\nabla_{\theta}\ell_{n}(\eta)\Bigr\|\leq C_{2}\,\delta_{n}\,$, (25)
where $\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ and $\nabla_{\theta}\ell_{n}(\eta)$ are the $\theta$}-score functions of $\widetilde{\ell}_{n}(\eta)$ and $\ell_{n}(\eta)$, respectively.
The proof of Lemma 8 follows directly from classical approximation results for Riemann sums. Under Assumptions A1–A4, $g(t,{\bf x})$ is uniformly Hölder continuous in $t$, uniformly over $\bf x$} and over $\eta\in\mathcal{N}_{n}$. Because the hazard is Lipschitz in $g(t,{\bf x})$, the integrand $Y_{i}(t)\exp\{g(t,{\bf X}_{i})+\mbox{${\theta}$}^{\top}{\bf Z}_{i}\}$ is Lipschitz on each interval where $Y_{i}(t)=1$. Therefore the Riemann approximation error is $O(\delta_{n})$. Applying this bound uniformly over all observations and all $\eta\in\mathcal{N}_{n}$ yields inequalities (24) and (25). Because the argument involves no stochastic elements and is fully deterministic, we omit the proof.
Lemma 9.
Let $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$, and $\dot{S}_{n}(\eta)=\nabla_{\theta}S_{n}(\eta)$ is the Hessian of the empirical log–likelihood with respect to $\theta$}. For a fixed $M>0$, define the shrinking neighbourhood $\mathcal{N}_{n}=\bigl\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\bigr\}$, $\tau_{n}=\gamma_{n}\log^{2}n$. Assume A1A4. Then,
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|\;\xrightarrow{P}\;0\,$,
where $I(\mbox{${\theta}$}_{o})$ is the efficient information matrix defined in Theorem 3, and here $\|\cdot\|$ may be any fixed matrix norm; for concreteness, and because it provides the appropriate control on the inverse Hessian, we take it to be the operator norm induced by the Euclidean norm.
Proof of Lemma 9.
For a single observation $(T,\Delta,{\bf X},{\bf Z})$, the per–observation log–likelihood is
$\ell(\mbox{${\theta}$},g)=\Delta\{\mbox{${\theta}$}^{\top}{\bf Z}+g(T,{\bf X})\}-\int_{0}^{\tau}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,.$
The corresponding $\theta$}-score is
$s(\mbox{${\theta}$},g)=\nabla_{\theta}\ell(\mbox{${\theta}$},g)=\Delta{\bf Z}-\int_{0}^{\tau}{\bf Z}\,Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,$,
and the $\theta$}-Hessian is
$\dot{s}(\mbox{${\theta}$},g)=\nabla_{\theta}s(\mbox{${\theta}$},g)=-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\,.$
Therefore,
$\dot{S}_{n}(\eta)=\mathbb{P}_{n}\dot{s}(\mbox{${\theta}$},g)$
and
$\mathbb{P}\,\dot{s}(\mbox{${\theta}$},g)=E_{0}\!\left[-\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt\right],$
where $\mathbb{P}_{n}$ is the empirical measure and $\mathbb{P}=P_{0}$ denotes expectation under the true law.
Define the matrix-valued per-observation information integrand by
$J(\eta;{\bf W})=\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\,dt,$
where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$ so that $\dot{S}_{n}(\eta)=-\,\mathbb{P}_{n}J(\eta;\cdot)$ and $\mathbb{P}\dot{s}(\mbox{${\theta}$},g)=-\,\mathbb{P}J(\eta;\cdot)$.
By Assumptions A1A4, the covariates are bounded, $\|{\bf Z}\|\leq B_{{\bf Z}}$ a.s., and the hazards are uniformly bounded. Then, there exist $0<c_{h}<C_{h}<\infty$ such that
$c_{h}\;\leq\;\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\;\leq\;C_{h}$
for all $\eta$ in the sieve and all $(t,{\bf X},{\bf Z})$. Hence the integrand in $J(\eta;{\bf W})$ is uniformly bounded by
$\|{\bf Z}{\bf Z}^{\top}Y(t)\exp\{\mbox{${\theta}$}^{\top}{\bf Z}+g(t,{\bf X})\}\|\;\leq\;B_{{\bf Z}}^{2}C_{h}$
for all $\eta\in\mathcal{N}_{n}$ and all $\bf W$}. Thus the class $\mathcal{J}_{n}=\bigl\{J(\eta;\cdot):\eta\in\mathcal{N}_{n}\bigr\}$ has a finite envelope and inherits the same entropy bounds as the likelihood class used in Lemma 4. In particular, $\mathcal{J}_{n}$ is $\mathbb{P}$–Glivenko–Cantelli, so
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|\;\xrightarrow{P}\;0\,.$ (26)
Fix $\eta_{1}=(\mbox{${\theta}$}_{1},g_{1})$ and $\eta_{2}=(\mbox{${\theta}$}_{2},g_{2})$ in $\mathcal{N}_{n}$. For each observation $\bf W$},
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|$ $=\Biggl\|\int_{0}^{\tau}{\bf Z}{\bf Z}^{\top}Y(t)\Bigl[\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr]\,dt\Biggr\|$
$\leq\int_{0}^{\tau}\|{\bf Z}{\bf Z}^{\top}\|\,Y(t)\Bigl|\exp\{\chi_{\eta_{1}}(t,{\bf X},{\bf Z})\}-\exp\{\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\}\Bigr|\,dt\,.$
Using the mean-value theorem for the exponential and boundedness of $\bf Z$} and $g$, there exists a constant $C>0$, depending only on model constants, such that
$\Bigl|\exp\{\mbox{${\theta}$}_{1}^{\top}{\bf Z}+g_{1}(t,{\bf X})\}-\exp\{\mbox{${\theta}$}_{2}^{\top}{\bf Z}+g_{2}(t,{\bf X})\}\Bigr|\;\leq\;C\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,.$
Therefore,
$\bigl\|J(\eta_{1};{\bf W})-J(\eta_{2};{\bf W})\bigr\|\;\leq\;C^{\prime}\int_{0}^{\tau}Y(t)\Bigl\{\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\Bigr\}\,dt$
for some $C^{\prime}>0$. Taking expectations and using $\int_{0}^{\tau}Y(t)\,dt\leq\tau$,
$\bigl\|\mathbb{P}J(\eta_{1};\cdot)-\mathbb{P}J(\eta_{2};\cdot)\bigr\|$ $\leq C^{\prime}\tau\|\mbox{${\theta}$}_{1}-\mbox{${\theta}$}_{2}\|+C^{\prime}\,E_{0}\!\left\{\int_{0}^{\tau}Y(t)\,|g_{1}(t,{\bf X})-g_{2}(t,{\bf X})|\,dt\right\}.$
By the Hölder regularity and boundedness assumptions on $g$, together with the fact that $d(\eta_{1},\eta_{2})$ controls the $L^{2}$–distance between $\chi_{\eta_{1}}$ and $\chi_{\eta_{2}}$, one can bound the right-hand side by $C^{\prime\prime}d(\eta_{1},\eta_{2})$ for some $C^{\prime\prime}>0$. In particular, for all $\eta\in\mathcal{N}_{n}$,
$\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\leq\;C^{\prime\prime}d(\eta,\eta_{o})\;\leq\;C^{\prime\prime}M\tau_{n}\;\longrightarrow\;0\,$,
since $\tau_{n}\to 0$. Hence
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|\;\longrightarrow\;0\,.$ (27)
By Theorem 3, the efficient score for $\mbox{${\theta}$}_{o}$ is
$\ell^{*}_{\theta_{o}}(T,{\bf X},{\bf Z},\Delta)=\int_{0}^{\tau}\{{\bf Z}-{\bf g}^{*}(t,{\bf X})\}\,dM(t)\,$,
and the efficient information matrix is
$I(\mbox{${\theta}$}_{o})=E_{0}\bigl[\Delta\{{\bf Z}-{\bf g}^{*}(T,{\bf X})\}^{\otimes 2}\bigr]\,.$
By semiparametric information theory, the efficient score satisfies the usual information identity along the least favorable submodel. In particular, the efficient Fisher information equals minus the expectation of the derivative of the efficient score,
$\mathbb{P}\dot{s}(\mbox{${\theta}$}_{o},g_{o})=I(\mbox{${\theta}$}_{o})\,.$
Recalling that $\dot{s}(\mbox{${\theta}$}_{o},g_{o})=-J(\eta_{o};{\bf W})$, this gives
$\mathbb{P}J(\eta_{o};\cdot)=I(\mbox{${\theta}$}_{o})\,.$ (28)
Finally, from (26), (27), and (28),
$\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\dot{S}_{n}(\eta)+I(\mbox{${\theta}$}_{o})\bigr\|$ $=\sup_{\eta\in\mathcal{N}_{n}}\bigl\|-\mathbb{P}_{n}J(\eta;\cdot)+I(\mbox{${\theta}$}_{o})\bigr\|$
$\leq\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}_{n}J(\eta;\cdot)-\mathbb{P}J(\eta;\cdot)\bigr\|+\sup_{\eta\in\mathcal{N}_{n}}\bigl\|\mathbb{P}J(\eta;\cdot)-\mathbb{P}J(\eta_{o};\cdot)\bigr\|$
$\xrightarrow{P}0+0\;=\;0.$
This proves the lemma. ∎
See 5
Proof of Theorem 5.
Recall that $\ell_{n}(\eta)$ and $\widetilde{\ell}_{n}(\eta)$ denote the exact and numerically integrated log-likelihoods, respectively, and $S_{n}(\eta)=\nabla_{\theta}\ell_{n}(\eta)$ and $\widetilde{S}_{n}(\eta)=\nabla_{\theta}\widetilde{\ell}_{n}(\eta)$ their $\theta$}-scores. By Lemma 8, for all $\eta$ in the shrinking neighborhood $\mathcal{N}_{n}=\{\eta:\ d(\eta,\eta_{o})\leq M\tau_{n}\}$ for some fixed $M\geq 1$, we have the uniform bounds
$\sup_{\eta\in\mathcal{N}_{n}}|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)|\;\leq\;C_{1}\delta_{n}\,$,
and
$\sup_{\eta\in\mathcal{N}_{n}}\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\|\;\leq\;C_{2}\delta_{n}\,.$
Since $\eta_{o}\in\mathcal{N}_{n}$ for all $n$ large enough, subtracting the values at $\eta_{o}$ gives
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl|\ell_{n}(\eta)-\widetilde{\ell}_{n}(\eta)-\{\ell_{n}(\eta_{o})-\widetilde{\ell}_{n}(\eta_{o})\}\bigr|=O_{p}(\delta_{n})\,$, (29)
and
$\sup_{d(\eta,\eta_{o})\leq M\tau_{n}}\bigl\|\widetilde{S}_{n}(\eta)-S_{n}(\eta)\bigr\|=O_{p}(\delta_{n})\,.$ (30)
These bounds reflect that the per-subject approximation error is $O(\delta_{n})$ and the network outputs are uniformly bounded because the sieve $\mathcal{G}(K,s,\boldsymbol{p},D)$ is constructed with bounded weights.
By (29) and Theorem 1, on the ball $\{d(\eta,\eta_{o})\leq\tau_{n}\}$ we have
$\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\widetilde{\ell}_{n}-\ell_{0})(\eta)-(\widetilde{\ell}_{n}-\ell_{0})(\eta_{o})\bigr|$ $\leq\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|(\ell_{n}-\ell_{0})(\eta)-(\ell_{n}-\ell_{0})(\eta_{o})\bigr|$
$\qquad+\sup_{d(\eta,\eta_{o})\leq\tau_{n}}\bigl|\widetilde{\ell}_{n}(\eta)-\ell_{n}(\eta)-\{\widetilde{\ell}_{n}(\eta_{o})-\ell_{n}(\eta_{o})\}\bigr|$
$=O_{p}(\tau_{n}^{2})+O_{p}(\delta_{n})=O_{p}(\tau_{n}^{2}),$
[…]
By the score equations, $S_{n}(\widehat{\eta})={\bf 0}$ and $\widetilde{S}_{n}(\widetilde{\eta})={\bf 0}$. Therefore,
[…]
Since $d(\widehat{\eta},\eta_{o})=O_{p}(\tau_{n})$ and $d(\widetilde{\eta},\eta_{o})=O_{p}(\tau_{n})$, there exists a constant $C>0$ such that, with probability tending to one, $\{d(\widehat{\eta},\eta_{o})\leq C\tau_{n},\;d(\widetilde{\eta},\eta_{o})\leq C\tau_{n}\}\subseteq\mathcal{N}_{n}$, where $M\geq C$. On this event, since the two estimators satisfy the score equations
$S_{n}(\widehat{\eta})={\bf 0}\,,\qquad\widetilde{S}_{n}(\widetilde{\eta})={\bf 0}\,$,
we also get
[…]
where $R_{n}(\eta)=\widetilde{S}_{n}(\eta)-S_{n}(\eta)$. Subtracting $S_{n}(\widehat{\eta})=0$ yields
[…]
In addition,
$S_{n}(\widetilde{\eta})-S_{n}(\widehat{\eta})=\underbrace{S_{n}(\widetilde{g},\widetilde{}\mbox{${\theta}$})-S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})}_{A}+\underbrace{S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})}_{B}.$
and (31) becomes
[…]
By the mean value theorem in $\theta$}, there exists $\mbox{${\theta}$}^{\dagger}$ on the segment between $\widehat{}\mbox{${\theta}$}$ and $\widetilde{}\mbox{${\theta}$}$ such that
$A=\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\,(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\,.$
By Lemma 9, $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})\xrightarrow{P}-I(\mbox{${\theta}$}_{o})$ and $\|\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}\|=O_{p}(1)$. Substituting this into (31), we obtain
[…]
and
[…]
Lemma 8 gives uniformly over $\mathcal{N}_{n}$,
$\|R_{n}(\eta)\|\leq C\delta_{n},$
so
$\sqrt{n}\,R_{n}(\widetilde{\eta})=O_{p}(\sqrt{n}\delta_{n})=o_{p}(1),$
because $\sqrt{n}\delta_{n}\to 0$ by assumption.
We decompose
$B=S_{n}(\widetilde{g},\widehat{}\mbox{${\theta}$})-S_{n}(\widehat{g},\widehat{}\mbox{${\theta}$})=\{\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\}+\{(\mathbb{P}_{n}-\mathbb{P})S(\widetilde{g},\widehat{}\mbox{${\theta}$})-(\mathbb{P}_{n}-\mathbb{P})S(\widehat{g},\widehat{}\mbox{${\theta}$})\},$
where $S(\eta;{\bf W})$ is the per–observation score in $\theta$} and $S_{n}(\eta)=\mathbb{P}_{n}S(\eta;\cdot)$. The second–order expansion of the population score (see the proof of Theorem 4) yields
$\|\mathbb{P}S(\widetilde{g},\widehat{}\mbox{${\theta}$})-\mathbb{P}S(\widehat{g},\widehat{}\mbox{${\theta}$})\|\lesssim\|\widetilde{g}-g_{o}\|_{L^{2}(\Delta)}^{2}+\|\widehat{g}-g_{o}\|_{L^{2}(\Delta)}^{2}=O_{p}(\tau_{n}^{2})\,.$
Moreover, Lemma 4 applied to the localized score class $\{S(g,\widehat{}\mbox{${\theta}$};\cdot):d((g,\widehat{}\mbox{${\theta}$}),\eta_{o})\leq C\tau_{n}\}$ implies
$\sup_{d((g,\widehat{\theta}),\eta_{o})\leq C\tau_{n}}|(\mathbb{P}_{n}-\mathbb{P})S(g,\widehat{}\mbox{${\theta}$};\cdot)|=O_{p}(\tau_{n}^{2})\,$,
and since both $(\widetilde{g},\widehat{}\mbox{${\theta}$})$ and $(\widehat{g},\widehat{}\mbox{${\theta}$})$ lie in this ball with probability 1o(1)1-o(1), we obtain
$\|B\|=O_{p}(\tau_{n}^{2}),\qquad\sqrt{n}B=o_{p}(1),$
because $\sqrt{n}\tau_{n}^{2}\to 0$. Using (32) together with $\dot{S}_{n}(\widetilde{g},\mbox{${\theta}$}^{\dagger})^{-1}=O_{p}(1)$, $\sqrt{n}B=o_{p}(1)$, and $\sqrt{n}R_{n}(\widetilde{\eta})=o_{p}(1)$, we obtain
$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})=o_{p}(1)\,.$
Finally, combining this with Theorem 4, which states that $\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})$ is asymptotically normal $N\bigl(\mathbf{0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)$, and using Slutsky’s theorem yields
$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\widehat{}\mbox{${\theta}$})\;\xRightarrow{d}\;N\bigl(\mathbf{0},I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,
as claimed. ∎